### v2

# Perceptions of the Fairness Impacts of Multiplicity in Machine Learning

Abstract.
Machine learning (ML) is increasingly used in high-stakes settings, yet multiplicity – the existence of multiple good models – means that some predictions are essentially arbitrary. ML researchers and philosophers posit that multiplicity poses a fairness risk, but no studies have investigated whether stakeholders agree. In this work, we conduct a survey to see how multiplicity impacts lay stakeholders’ – i.e., decision subjects’ – perceptions of ML fairness, and which approaches to address multiplicity they prefer. We investigate how these perceptions are modulated by task characteristics (e.g., stakes and uncertainty). Survey respondents think that multiplicity threatens the fairness of model outcomes, but not the appropriateness of using the model, even though existing work suggests the opposite. Participants are strongly against resolving multiplicity by using a single model (effectively ignoring multiplicity) or by randomizing the outcomes. Our results indicate that model developers should be intentional about dealing with multiplicity in order to maintain fairness.
1. Introduction
Machine learning (ML) is ubiquitous in high-stakes decision-making settings, e.g., social services (Chouldechova et al., 2018), criminal justice (Propublica, 2016), healthcare (Obermeyer et al., 2019), and hiring (Li et al., 2021). Due to the life-altering impacts of these decisions (for instance, whether someone can get hired at a job), it is important that the model making the decision is as accurate, fair, and accountable as possible. The conventional machine learning wisdom is that there exists a single best-performing model and that achieving other desiderata like fairness or robustness requires an accuracy tradeoff (Menon and Williamson, 2018). But the literature – both recent (D’Amour et al., 2022; Marx et al., 2020) and foundational (Breiman, 2001) – challenges that paradigm: there are typically many models that achieve similar predictive accuracy by different means (i.e., by making mistakes on different parts of the input distribution). This ambiguity leads to predictive multiplicity, which is when two equally good models assign different predictions to a given sample, as illustrated by fig. 1. In socially-salient prediction settings, the set of (approximately) equally good models is often very large (Marx et al., 2020; Rudin et al., 2024; Xin et al., 2022), and a large fraction of samples receive conflicting outcomes under this multiplicity (Marx et al., 2020).
Work in computer science and philosophy suggests that multiplicity poses a fairness risk (Black et al., 2022b; Cooper et al., 2024; Creel and Hellman, 2022). Fairness has multiple components, including distributional fairness, which considers whether the outcomes are fair, and procedural fairness, which considers whether the decision was made in a fair way, regardless of outcomes (Colquitt, 2001). The existing work points out that multiplicity can violate procedural fairness, because choosing one model over the alternatives results in arbitrary decisions (Black et al., 2022b; Cooper et al., 2024; Creel and Hellman, 2022). This fairness violation is especially salient when models operate in high-stakes settings where the justifiability of a negative decision is important. For instance, suppose a company uses an ML model to determine which job applicants to interview, but that there is a different model that would do equally as well at selecting promising candidates (e.g., and have similar accuracy on the available validation data). A candidate who is not interviewed according to , but who would have been interviewed if were used, may feel that procedural fairness is violated if the company did not even consider using . An additional concern is that when models are deployed at scale, multiplicity can result in systematic arbitrariness that locks people out of opportunities (Creel and Hellman, 2022). For instance, if all companies contract with the same software vendor and each uses model to determine who to interview, may get no job interviews even though the existence of suggests they may be qualified.
While the existing literature makes a strong argument for the fairness risks of multiplicity, no research has interrogated whether lay stakeholders believe that multiplicity impacts the fairness of machine learning models. Likewise, no research has asked how stakeholders think decisions should be made when a machine learning system contains multiplicity. More precisely, consider the same job application scenario where two models and have similar performance, yet disagree on whether to interview a job candidate . The standard machine learning approach is to choose a single well-performing model, for instance, selecting without even realizing that exists. Philosophers suggest using randomization – e.g., flipping a coin to determine whether to interview – as the fairest way to deal with multiplicity (Creel and Hellman, 2022; Jain et al., 2024). Computer scientists suggest switching to more sophisticated model architectures to reduce multiplicity (e.g., ensemble methods as a way to reduce variance) (Black et al., 2022a; Cooper et al., 2024; Roth et al., 2023). Lay stakeholders will not necessarily like those approaches, as research shows that different types of stakeholders (like decision makers and decision subjects) can have different perceptions of what matters for fairness in ML (Lee and Rich, 2021; Scott et al., 2021; Zhang et al., 2023). Critical scholars argue that considering stakeholder preferences is important (Bondi et al., 2021; Selbst et al., 2019) and this belief underlies frameworks like participatory ML (Feffer et al., 2023) and value-sensitive design (Umbrello and Van de Poel, 2021). Furthermore, companies deploying ML may care about public opinions of fairness to maintain high levels of trust in their product (Woodruff et al., 2018).
We address the gaps in understanding stakeholders’ views on multiplicity by conducting a survey of crowdworkers (i.e., laypeople who are decision subjects as ML is increasingly deployed in everyday settings). Our survey aims to determine (1) how the presence of multiplicity impacts the stakeholders’ assessment of ML models’ fairness and (2) which multiplicity resolution approaches they prefer. We consider the following research questions:
The image contains 3 sub-figures, labeled A, B, and C. Image A has a cluster of data points (red circles and blue triangles) that are not perfectly linearly separable. The image shows two linear classifiers that each misclassify 1 point (but a different point for each). The points that are misclassified by a model are circled in yellow. Image B is formatted as a table where headers are pictures of animals (Column 1 Cat, Column 2 Dog, Column 3 cat). Each row contains the prediction that an AI system makes on each image (predicting whether it is a cat or dog). The first row (AI 1) is correct on examples 1 and 3, while the second row (AI 2) is correct on examples 2 and 3. Text in the image indicates that both AIs are correct on 2 out of 3 examples. Image C is a venn diagram. There is a large circle labeled All Reviews. Within this circle, there are 3 smaller circles arranged along an invisible diagonal line. The middle circle (green) is labeled Reviews Deleted by Human Moderator. The other circles are labeled Reviews deleted by the original model (orange) and Reviews deleted by the new model (pink). The orange and pink circles overlap about 50 percent of their area, while each overlaps with the green circle by about 75 percent of their area.
RQ1: Does being alerted to multiplicity in a machine learning system affect participants’ perceptions of fairness of the system?
RQ2: What techniques to resolve multiplicity do participants view as most appropriate and fair?
RQ3: How do task characteristics (stakes, uncertainty, and punishment or reward framing) impact RQ1 and RQ2?
To address these questions, we want to evaluate fairness perceptions for a variety of tasks. However, some of the factors we care about for RQ3 (namely, task stakes and uncertainty) may be subjective, so we did not want to rely solely on our own classification of tasks. Instead, we conducted a preliminary study with participants, which, to the best of our knowledge, is the first study to get lay stakeholders’ opinions on both tasks’ perceived stakes and uncertainty (Castelo et al. (Castelo et al., 2019) collect opinions on task uncertainty, but not stakes). Based on the results of this survey and our own assessment of task framing (reward vs. punishment), we selected 8 diverse tasks to use in the main study.
Our main study begins with an educational component describing model multiplicity. After the participant passes a comprehension check, we assign them to a scenario and describe a machine learning model at a high level before asking participants for an initial fairness assessment, consisting of a series of Likert scale questions. Then, we describe how model multiplicity impacts the situation and ask for updated answers to the same Likert questions (RQ1). Finally, we use an open-ended question and choice-based analysis to learn what multiplicity resolution techniques participants view as the most fair (RQ2).
Our results indicate that stakeholders do not perceive multiplicity as an overall fairness risk, counter to what the literature suggests. However, this result is nuanced as participants are more likely to express that the model outcomes might be unfair, rather than that the model may be inappropriate to use at all. We also find that participants have clear preferences for how multiplicity is resolved, namely by deferring to a human expert for situations where multiplicity impacts the predictions (RQ2). These preferences vary depending on the specific task, with task stakes and task framing significantly impacting resolution preferences (RQ3). Despite philosophical arguments for randomization as a multiplicity solution (Creel and Hellman, 2022; Jain et al., 2024), this was the least preferred option in our experiment.
These results have two important implications: first, the way that multiplicity is discussed in the existing literature does not align with the lay stakeholders’ perceptions. Instead, we hypothesize that other concerns about machine learning (such as recourse and explainability) may dominate participants’ fairness assessments. Second, even though overall fairness perceptions do not shift significantly after learning about multiplicity, the strong sentiments that participants have regarding multiplicity resolution suggest that ML developers must be deliberate in determining how models behave in the presence of multiplicity. In particular, simply choosing a single well-performing model and not exploring the impacts of multiplicity is viewed as unfair.
2. Background of Fairness and Multiplicity
We overview fairness in the context of machine learning (section 2.1), define multiplicity and its associated fairness risks (section 2.2), and describe various ways to resolve multiplicity (section 2.3).
2.1. Fairness in machine learning
Colquitt and Rodell propose several components of fairness including distributional, procedural, and informational fairness (Colquitt and Rodell, 2015). Distributional fairness asks whether a model’s outcomes are fair, irrespective of how the decisions were made. For instance, group fairness metrics such as equalized odds (Dwork et al., 2012; Hardt et al., 2016) ask whether different demographic groups are treated equally in aggregate. Procedural fairness, by contrast, asks whether a decision was made in a fair way, regardless of what the final outcome is. Procedural fairness has many facets, including counterfactual fairness (Kusner et al., 2017) (how a test sample needs to change to get a different prediction) and the choice of features that are included in a model. For example, in a task such as hiring, procedural fairness may be violated if race is a decisive factor (i.e., changing an applicant’s race changes whether they are hired) or even if race is included as a model input at all. Informational fairness requires that details about the decision making process be available, and has been explored in ML through the use of explanations, e.g. (Binns et al., 2018; Dodge et al., 2019; Schoeffer et al., 2022; Yurrita et al., 2023). In this paper, we consider perceptions of procedural fairness.
The fairness of a ML model can be measured based on its adherence to some fixed standard (e.g., a group fairness metric), or in a human-centered way through perceived fairness, as in this study. Perceived fairness is a common way to evaluate ML systems (Starke et al., 2022), perhaps because settling on a formal fairness definition is challenging. Research has established that perceived fairness of a model correlates with trust of that model (Shin et al., 2020; Shin and Park, 2019; Woodruff et al., 2018), and suggests that low perceived fairness corresponds to negative consequences (e.g., reputational harm) for the entity deploying ML (Starke et al., 2022).
2.2. Multiplicity in machine learning and fairness concerns
Predictive multiplicity (Marx et al., 2020), which we simply call multiplicity for brevity, occurs when there are multiple ML models that are equally well-suited for a task, yet give different individual-level predictions. Formally, suppose we have a dataset , a hypothesis class (e.g., all possible decision trees with depth ) and some loss function (e.g., the 0-1 loss). A standard machine learning pipeline aims to find that minimizes (e.g., the total number of misclassifications on ). This solution is typically found via an approximate optimization process such as gradient descent, and yields a model that hopefully has close-to-optimal loss, i.e., has . But sometimes, the optimization problem is underconstrained and there is another model such that . (Intuitively, there may be multiple global minima.) If there is some input with we say that the models exhibit (predictive) multiplicity.
Prevalence of multiplicity
Computing the fraction of inputs whose predictions are affected by multiplicity is generally intractable, so existing approaches quantify multiplicity by limiting their analysis to specific model types, approximating the impacts of multiplicity, and/or using brute force. Estimates on the COMPAS dataset, which is used to predict risk of recidivism in defendants, indicate that multiplicity affects 44% of test inputs when using linear models (Marx et al., 2020) and 69.4% of test inputs when using decision trees (Cooper et al., 2024). For probabilistic models that compute risk scores, model outputs vary by on 82-100% of samples across three tabular datasets (Watson-Daniels et al., 2023). These levels of multiplicity indicate that many predictions are a function of the specific model that is chosen, rather than purely a function of the data.
Fairness risks of multiplicity
Fairness risks of multiplicity include harms to an individual, society or a group of individuals, and the entity who is deploying the model. Fairness risks to individuals include a lack of justifiability for a negative decision (Black et al., 2022b) and systemic loss of opportunity when models containing multiplicity are deployed at scale (Creel and Hellman, 2022).
Example 2.1 (Lack of justifiability).
Black et al (Black et al., 2022b) point out that US credit laws like the Equal Credit Opportunity Act aim to reduce arbitrariness and increase equity in credit decisions, such as by requiring lenders to provide justification for denied loan requests. However, if another multiple equally-accurate model approves a loan request, it may be more challenging to justify why the bank needed to use the model that denies the loan.
Example 2.2 (Systemic loss of opportunity).
Creel and Hellman (Creel and Hellman, 2022) discuss how individual employers may have idiosyncratic hiring criteria, but that the harm from this to job applicants is minimal, as they can apply to many jobs. Since each employer has their own potentially arbitrary criteria (in other words, multiplicity exists in the process of choosing who to hire), the job applicant will eventually be hired by someone. However, if every company uses the same hiring system, then certain job applicants will not be able to be hired anywhere, causing them significant harm. This setup can happen, for instance, if every company contracts with the same third-party software to screen resumes.
Society or groups of individuals may be adversely impacted by multiplicity if a sub-optimal model, with regards to fairness, is selected from the set of equivalently-accurate models. The entity using the model is also at risk, first for potential lawsuits or reputational harm if their decisions are not justifiable (Black et al., 2022b). Second, if a single model is deployed at scale when multiplicity exists, this standardization can provide a worse set of outcomes.
Example 2.3 (Worse outcomes for the decision-maker).
Kleinberg et al. (Kleinberg and Raghavan, 2021) show that companies who all use the same recruiting software may end up with lower-quality new hires than if each company used its own model, an outcome that is harmful for the company itself and not just potential job applicants.
However, no work that we are aware of has investigated the fairness risks of multiplicity from the perspective of stakeholders. Previous work shows that formal fairness definitions often do not line up with stakeholders’ fairness perceptions (Srivastava et al., 2019), so, it seems likely that experts and lay stakeholders may have distinct insights about whether multiplicity is (un)fair.
2.3. “Solving” multiplicity
A natural question, after finding evidence of multiplicity, is how to go about making decisions. In this paper, we consider the fairness of six different multiplicity-resolution techniques (see table 1). The standard approach in machine learning is to ignore111Here, and throughout the rest of the paper, we use this font to denote the names of the multiplicity-resolution techniques. multiplicity; that is, to find a single well-performing model and use it, without considering whether multiplicity exists. Of work that focuses on multiplicity, three schools of thought exist on how to resolve it: reducing multiplicity through algorithmic interventions, standardizing the procedural response to multiplicity, and using multiplicity as an opportunity to optimize for second-order criteria like fairness or robustness.
Reducing multiplicity
Refining the modeling technique can reduce multiplicity, such as by using cascading classifiers (Roth et al., 2023), bootstrapping (Cooper et al., 2024), or selective ensembles that abstain if the confidence is not high enough (Black et al., 2022a). Similar ensembling approaches have long been used to improve accuracy in ML applications (Sagi and Rokach, 2018). The difference when using ensembling to combat multiplicity is that multiplicity is often discussed in the context of social prediction tasks where an objective ground truth is elusive or even does not exist. Thus, using ensembling to get rid of multiplicity often does not yield significantly higher accuracy than using a non-ensembled model. It does, however, mask that multiplicity exists, making it harder to challenge the validity of an algorithm. We refer to ensembling and other approaches that aim to reduce multiplicity through modeling choices as fancy, i.e., training a fancier (more complex) model.
| Method name | When the two models disagree… |
| default-bad | Default to the undesirable outcome |
| default-good | Default to the desirable outcome |
| fancy | Train a more complex model |
| and use its output | |
| human | Let a human expert decide |
| ignore | Use the original model’s decision |
| random | Choose the outcome randomly |
Dealing with multiplicity procedurally
ML researchers have suggested abstaining, and, presumably, deferring to a human expert (denoted in this paper as human), when a decision is subject to multiplicity (Black et al., 2022a). However, the design of “human-in-the-loop systems” is not straightforward (Wu et al., 2022), and may interact with other criteria we are trying to measure, e.g., algorithm aversion (Fogliato et al., 2022). The idea of abstaining and deferring to a human decision-maker approach parallels a long line of work in robust machine learning that similarly suggests abstaining on decisions when robustness cannot be verified (Bartlett and Wegkamp, 2008; Cohen et al., 2019; Herbei and Wegkamp, 2006; Laidlaw and Feizi, 2019).222Note that non-robustness constitutes a form of multiplicity, as it implies there is another reasonable model that outputs an alternate prediction for the test sample. Within the ML literature, robustness is typically discussed for deep networks and/or text and image data, while multiplicity research largely focuses on social prediction tasks using tabular data. However, these works rarely present a plan for what abstaining looks like in practice: often, abstaining on a decision results in not taking action, thereby defaulting to the negative label (for instance, abstaining on a loan application is effectively denying it). We consider abstaining and deferring to a default option in our survey. However, attitudes may vary depending on whether the outcome is desirable or not, so we use default-good and default-bad to represent the outcomes that are good and bad, respectively, from the perspective of the decision subject.
An alternate procedural approach is random, i.e., randomizing over the set of equally-good models (Creel and Hellman, 2022; Jain et al., 2024). ML in social prediction tasks often serves to distribute a limited resource, e.g., admissions to a college with a fixed enrollment limit. Lotteries are typically viewed as fair system for dividing limited resources (Broome, 1990; Sher, 1980). However, there are technical challenges in implementing randomization effectively (Jain et al., 2024).
Optimizing for second-order criteria
Traditional ML practice assumes that there is a trade-off between model accuracy and fairness, as defined by group-level metrics such as equalized odds or statistical parity (Kamiran and Calders, 2012; Menon and Williamson, 2018). Recent research to explore the whole set of equally-good models challenges this paradigm, showing that accuracy and other criteria are not in tension (Rudin et al., 2024; Xin et al., 2022). But even if we can use multiplicity to find fairer (or more robust) models, there are still multiple mutually incompatible fairness (or robustness) definitions. We do not include optimizing for fairness as one of the possible multiplicity resolution techniques because the task of choosing the fairest model is typically not well-defined (Chouldechova, 2017; Kleinberg et al., 2017).
3. Related Work and Hypotheses
Fairness perceptions of machine learning algorithms have been studied extensively, as evidenced by multiple survey papers on the topic (Narayanan et al., 2024; Starke et al., 2022). Here, we will focus on dimensions most related to our research questions.
Research in this area investigates distributional fairness (e.g., matching lay stakeholder’s fairness notions to formal definitions (Saxena et al., 2019; Srivastava et al., 2019)), procedural fairness (e.g., considering what features to include (Grgic-Hlaca et al., 2018)), and informational fairness (e.g., the impact of explanations (Schoeffer et al., 2022; Yurrita et al., 2023)). In this work, we focus on perceptions of procedural fairness, presenting multiplicity as a hurdle that must be dealt with procedurally. The fairness perceptions of ML models are often considered relative to a baseline of human decision making, and can either increase from algorithmic appreciation (Logg et al., 2019), or decrease from algorithmic aversion (Dietvorst et al., 2015). Existing work establishes that fairness perceptions of ML models vary based on task characteristics, including stakes (Hannan et al., 2021; Saragih and Morrison, 2022; Yurrita et al., 2023), uncertainty (Salimzadeh et al., 2024), complexity (Salimzadeh et al., 2023), and framing (reward vs. punishment) (Hannan et al., 2021).
In the following subsections, we describe our hypotheses about how multiplicity impacts perceived fairness (H-H) and which multiplicity resolution techniques stakeholders prefer (H-H). See table 3 for a summary of all hypotheses.
3.1. Algorithm aversion
Algorithm aversion is when people prefer human decision-makers, even when the available AI system generally outperforms humans (Dietvorst et al., 2015). It may be triggered when people see a model make a mistake, even if they would have been forgiving of a human making the same mistake (Dietvorst et al., 2015), or when the model’s performance is worse than stated (Yin et al., 2019). In settings where a correct answer can take many forms (e.g., generative AI), seeing multiple conflicting output lowers trust in a model (Lee et al., 2024). Algorithm aversion also tends to be higher for subjective tasks, a characterization that is malleable depending on how the task is described (Castelo et al., 2019). We expect that witnessing model multiplicity will increase algorithm aversion for three reasons: first, the existence of two good (but different) models means that the models make mistakes, similar to what was observed for conflicting generative AI responses (Lee et al., 2024). Second, the existence of multiple good models indicates that the decisions these models make are subjective. In particular, multiplicity means that choosing a model is not a purely objective task (i.e., optimizing accuracy), but rather a subjective choice of choosing between multiple models with near-equivalent accuracy. Finally, existing work assumes (without consulting stakeholders) that multiplicity threatens fairness (section 2.2).
H: Stakeholders will view machine learning systems as less fair once they are made aware of multiplicity.
Between algorithm aversion and prior work showing preference for human over AI arbitrators (Sela, 2018), we think participants will prefer human over other resolution techniques (recall that human and the other approaches are defined in table 1). We think that – counter to what philosophical work suggests – random will be viewed as the least fair approach. Explanations are important to fairness (Schoeffer et al., 2022; Yurrita et al., 2023), so we think the sense of unpredictability from randomized decisions dominate stakeholders’ reactions to random.
H: Stakeholders will have significant preferences between different multiplicity solutions. In particular, they will view human as the fairest multiplicity solution and random as the least fair solution.
3.2. Task stakes
Competing with algorithm aversion is algorithm appreciation, when people over-rely on algorithmic decisions (Logg et al., 2019). There is evidence that algorithm appreciation occurs in high-stakes tasks (Araujo et al., 2020; Saragih and Morrison, 2022), and stakeholders rate accuracy as more important in high-stakes settings (Srivastava et al., 2019). Thus, we think that fancy (i.e., building a more complex model) may be preferred for high-stakes tasks because it may be perceived as being more accurate. However, the importance of accuracy may also trigger algorithm aversion and increase preference for human. Fairness is more important to stakeholders in high-stakes settings than low-stake ones (Jakesch et al., 2022). We suspect that higher-stakes tasks are more emotionally charged. People view emotional decision making as a strength of humans (Helberger et al., 2020), and can have negative emotional responses to AI decision-making (Lee, 2018).
H: The presence of multiplicity will have a larger impact on fairness perceptions in high-stakes settings.
H: Participants’ preferred multiplicity resolution technique will be moderated by task stakes. In particular, high-stakes tasks will correspond to stronger preferences for human and fancy, and low-stakes tasks will correspond to stronger preferences for ignore and random.
3.3. Task framing
Harm that is caused by action, rather than omission, causes harsher fairness judgments (Cushman et al., 2006). Tasks that are punitive commit active harm in the event of a mistake, whereas reward-based tasks cause harm when they make an omission. Thus, we think that punitive tasks will have greater decreases in fairness perceptions in the presence of multiplicity. Indeed, others have found that when people adopt critical mindsets, they are more likely to find issues with a model (Hannan et al., 2021). Algorithms are seen as more rational than humans (Logg et al., 2019), and we suspect that rationality will be more important in punitive settings where mistakes are judged more harshly. Conversely, there are positive connotations towards human accuracy (Lee, 2018), so being primed to think about reward settings may influence participants to select human.
H: The presence of multiplicity will have a larger impact on fairness perceptions in punitive settings.
H: Preferences about how multiplicity is resolved will be moderated by task framing. In particular, for punitive tasks, fancy will be more preferred and for reward-based tasks, human and default-good will be more preferred.
3.4. Task uncertainty
High-uncertainty tasks typically do not have a single objectively-right answer, such as predicting future human behavior (e.g., job performance) or matters of taste (e.g., movie recommendation). Low-uncertainty tasks typically have an objectively-correct answer (e.g., detecting spam emails). Algorithm appreciation is known to be stronger in objective (i.e., low-uncertainty) tasks (Bankins et al., 2022; Logg et al., 2019), aligning with beliefs that algorithms are not good at subjective (i.e., high-uncertainty) tasks (Castelo et al., 2019). There is more algorithm aversion and preference for human decision makers in high-uncertainty settings (Dietvorst and Bharti, 2020). When a task has lower uncertainty, it may be possible to have overall higher accuracy, which increases algorithm appreciation (Saragih and Morrison, 2022). Intuitively, people may find it easier to rationalize that there are multiple good solutions to a high-uncertainty task, resulting in lower impact to the perceived fairness.
H: The presence of multiplicity will have a larger impact on fairness perceptions for low-uncertainty tasks.
H: Preferences of how to resolve multiplicity will be moderated by task uncertainty. In particular, for low-uncertainty tasks stakeholders will see fancy as more fair, while for high-uncertainty tasks, they will see random as more fair.
4. Method
Figure 2 shows an overview of our study design, which was approved by our institution’s IRB board. In this section, we first describe our preliminary study (section 4.1), which informed the task selection for our main study. Then, we describe our pre-registered study333https://osf.io/fv3bt design (section 4.2) and data analysis plan (section 4.4).
Visual overview of our study design. There are two rows. First row text is Preliminary study (n=199). Goal: determine task stakes and uncertainty to select tasks for the main study. There are 18 tasks total and each person sees 4. We ask How high-stakes is the task (Likert), How much uncertainty is there in the task outcomes (Likert), and we ask the participants to rank the 4 tasks they see on stakes and on uncertainty. The second row says Main Study (n=358). Goal: determine how multiplicity impacts fairness (Part1) and how participants think multiplicity should be resolved (Part 2). There are 8 tasks total and each person sees 1. The remainder of this row is subdividing into Part 1 and Part 2. In part 1, the flow is Comprehension (education on what multiplicity is) to Initial fairness perceptions (Likert) to Critical thinking (reveal that multiplicity exists, open-ended response), to Final fairness perceptions (Likert). In Part 2, it says Choice Based Analysis (we ask do you prefer option A or Option B). We randomly populate a table with two rows and two columns. The first column is labelled option A and the second column is labelled option B. The first row is How to resolve multiplicity and the second row is Cost. The example shown in the image has Option A= Choose randomly and None and Option B = Human employee and 5 dollars.
4.1. Task selection (Preliminary study)
Our main study asks participants about tasks that differ on three between-subjects axes: stakes, uncertainty, and framing. Framing (i.e., whether a positive model output is a reward or a punishment) is objective; however, individuals may differ on how high-stakes or uncertain they think a task is. We conduct a preliminary study () to solicit perceptions of task stakes and uncertainty. Our goal is to select one task for each of the eight categories [punishment, reward] [low uncertainty, high uncertainty] [low stakes, high stakes].
Study design
We collect 18 scenarios through brainstorming and borrowing from previous studies (Binns et al., 2018; Castelo et al., 2019; Kasinidou et al., 2021). We randomly assign each participant to see 4 scenarios to limit participant fatigue when completing the study. For each scenario, participants answer a series of 5-point Likert-scale questions relating to task clarity (2 questions), task stakes (3 questions), and task uncertainty (2 questions). Then, participants rank the four tasks according to stakes and uncertainty levels. We use Prolific to find respondents (US-based adults with a 97%+ approval rate). We compensate participants at Prolific’s recommended rate ($12/hour) based on the estimated survey completion time.
Data analysis
Our survey yields four pieces of data: stakes rating, uncertainty rating, stakes ranking, and uncertainty ranking. We convert all Likert ratings to a scale from -2 (strongly disagree) to 2 (strongly agree). For the rating data, we average the three questions relating to task stakes into a single stakes rating quantity. We elect to use one of the task uncertainty questions alone, due to better data quality. To recover the underlying stakes and uncertainty rankings given the partial ranking information from each participant, we use assume a Plackett-Luce distribution (Luce, 1977; Plackett, 1975) and use the I-LSR (iterative Luce Spectral Rating) algorithm (Maystre and Grossglauser, 2015). The basic assumption behind Plackett-Luce is that choice rankings are transitive (i.e., if A¿B and B¿C then A¿C), and the I-LSR algorithm uncovers the full underlying ranking. We combine the rating and ranking data into a single metric by scaling each set of data to have mean 0 and variance 1, and then taking their arithmetic mean.
Preliminary study results
We obtain responses from 201 participants and, after excluding responses that failed more than one attention check, use 199 responses in the analysis.
Task stakes and task uncertainty are not meaningfully correlated (=0.19, =0.42). We exclude three tasks that scored poorly on clarity metrics, leaving 15 tasks to choose from. There are trade-offs in terms of achieving either a wide range of scores on each axis or relative parity between tasks in different (stakes uncertainty) quadrants. That is, we can either select tasks to maximize the range of stakes and uncertainty scores, or we can select tasks with a narrower range of scores but with a better one-to-one comparison between tasks with different framing. We decide to maximize the range stakes and uncertainty scores since our analysis plan involves using numeric stakes and uncertainty labels rather than solely binary “low” and “high” categories. The tasks we select are described in table 2.
| Domain | Task | Stakes* | Uncertainty* | Framing |
| Healthcare | Access to a preventative health care plan | High | High | Reward |
| Business | Job probation based on employee monitoring | High | High | Punishment |
| Healthcare | Tumor detection from medical imaging | High | Low | Reward |
| Government | Tax audit | High | Low | Punishment |
| Travel | Airline seat upgrade | Low | High | Reward |
| Travel | Rebooking passengers on an over-full flight | Low | High | Punishment |
| Business | Customers get beta access to an app with discounts | Low | Low | Reward |
| Business | Removing online reviews that are spam or fake | Low | Low | Punishment |
4.2. Study design
The survey flow is depicted in Figure 2. The survey begins with education about multiplicity (with a comprehension question) and then has two main sections (“Part 1” and “Part 2”). Participants are randomly assigned to one task at the beginning of Part 1. Then we describe a scenario and solicit the participant’s fairness perceptions of the model on a Likert scale consisting of four seven-point questions. In Part 2, we present a series of binary choices between different ways to resolve multiplicity and the user selects their preferred option. We conclude by collecting demographic information from the participant. We provide further details on all parts of the survey below.
Details of multiplicity education
We describe to participants, at a high level, how AI models are built (e.g., “AI developers […] choose the model with the best performance”). Then, we reveal that there are sometimes multiple models that can perform equally well, but “make different mistakes”. We illustrate model multiplicity using a graphic similar to fig. 1B. We tell participants that this phenomenon, model multiplicity, is common. Then, participants have two chances to correctly answer the multiple-choice question “What does model multiplicity mean in the context of AI?” The correct answer is written using similar phrasing as in the descriptive text. We include full text of the multiplicity education section in the appendix.
Details of Part 1
In Part 1, we aim to address RQ1 and RQ3 by seeing how participants’ fairness perceptions of ML systems change when multiplicity is present. We first describe the scenario including what decision is being made, what factors the AI has access to, and how decisions will be operationalized (e.g., whether users will have recourse). We also say that the AI is validated and “agrees with human experts most of the time”. This description of accuracy is vague to avoid anchoring participants to specific numbers and because reporting an “accuracy” suggests a ground truth that does not exist for all scenarios we consider.
Next, we ask users to respond to four seven-point Likert scale questions to elicit fairness perceptions. Two questions focus on the model outcomes, e.g., whether the proposed model is likely to be able to make “good” decisions, and two questions focus on whether using the model is appropriate. Then, we describe how multiplicity impacts the situation, explaining that there is a second model that agrees with human experts equally often as the first model, but disagrees less often with the original model. We accompany the description with a graphic such as the one shown in Figure 1C. At this point, participants answer two critical-thinking questions: one is multiple-choice and comprehension based; the other is open-ended and asks how decisions should be made when the two models disagree.
The next page of the survey asks the participant to respond again to the same set of Likert questions. We refer to the first set of responses as the “pre” timestamp, and this set of responses as the “post” timestamp, because the latter are collected after the participant learns that multiplicity exists for the task. The participant does not have access to their previous responses at the “post” timestamp, because we do not want them to be anchored by their original responses (i.e., we want to allow for a change in their frame of reference (Howard et al., 1979)).
Details of Part 2
In part 2, we investigate RQ2 and RQ3 by seeing which multiplicity resolution techniques participants prefer. We provide a description of 5-6 ways (depending on the task) that multiplicity can be resolved, i.e., how to make decisions when the two models disagree (see table 1). For some tasks, default-good is not realistic due to logical constraints, e.g., airline seat upgrades likely cannot be given to every passenger selected by either model because there is a firm limit in the number of available first-class seats. In cases like this, we exclude that option. We also explain to participants that some options may be more expensive to implement, so might require a trade-off such as a monetary cost. The cost that participants see for each option is either “no extra cost” or a small scenario-specific cost that is meant to be an annoyance, but not a barrier for most people (e.g., airline tickets will cost $5 extra).
The participant then sees 8 randomly constructed binary choices between two randomly chosen multiplicity resolution techniques and associated costs, such as the example in fig. 2.
Demographic information
We collected demographic information from participants including gender, age, education level, and experience with AI. Experience with AI was collected through the question “Do you interact with artificial intelligence or automated decision-making systems as part of your work or school?”, with the choices “yes,” “no,” and “not currently, but in the past”.
4.3. Data collection
We collect data using Prolific. Our participation criteria is US-based adults (18+) who have at least a 95% approval rate. We plan to collect data from 360 participants, i.e., 45 per task. We estimated that 35-40 participants per task would be sufficient based on trends in pilot data; we rounded up to 45/task to have additional buffer in case we need to reject some samples. We follow Prolific’s guidance (Prolific, 2024) and exclude survey respondents who fail the initial comprehension check question twice, fail two or more attention check questions, or are statistical outliers in survey completion time (in the fast direction).
4.4. Statistical analyses
4.4.1. Part 1 analysis
Data transformation
We recode the Likert scale data from -3 (strongly disagree) to 3 (strongly agree) and compute the average across the four Likert scale questions.
Testing H-H
To test H-H, we fit a linear mixed model (LMM), which generalizes a linear model by accounting for both between- and within-subject factors. The formula we fit to the LMM predicts the transformed Likert data as a function of time, task stakes, task framing, task uncertainty, respondent ID (to account for individual differences), and the first-order interaction terms of time with each task characteristic. We use the lme4 package (Bates et al., 2015) in R to construct the LMM model and then use the car package (Fox and Weisberg, 2019) to conduct a type-III ANOVA. Using ANOVA procedures on Likert-scale data is controversial because Likert scales are not continuous and may not produce normally distributed data. Scholars have argued both for (Blanca Mena et al., 2017; Carifio and Perla, 2008; Norman, 2010) and against (Jamieson, 2004; Kuzon et al., 1996) applying ANOVA when its assumptions are violated. We decided that LMMs fit our needs best, so to minimize concerns about using ANOVA, we included a large population in each group (as required by the central limit theorem for normality of the data sample) and used the average of several Likert items rather than a single item (Carifio and Perla, 2008). The main alternative technique we considered is Aligned-Rank-Transform (ART) ANOVA (Wobbrock et al., 2011), a procedure that transforms ordinal (e.g., Likert-scale) data to fit the ANOVA assumptions. Ultimately we chose LMMs because ART-ANOVA does not allow for continuous factors. However, for readers who are interested, we include two additional analyses in the appendix: ART-ANOVA and a non-parametric test, the Mann-Whitney U-Test (Mann and Whitney, 1947).444To clarify, we preregistered the analysis with LMMs. Readers should view the results of the other tests as a partial multiverse analysis (Steegen et al., 2016).
To test H, we check whether the ANOVA p-value is significant, and if so, run a post-hoc contrast test to gauge directionality. To test H and H, we find the estimated marginal trends for how task stakes (H) and uncertainty (H) vary with time using emtrends in the emmeans package in R (Searle et al., 1980). Then, we conduct a pairwise contrast test on the result. To test H, we perform a pairwise contrast test on the estimated marginal means of the interaction term of time and task framing, again using emmeans.
Multiple hypothesis correction
In total, since we perform four post-hoc contrast tests on the Likert scale data, we will use the Holm-Bonferroni correction (Holm, 1979) with and to evaluate the hypotheses.
4.4.2. Part 2 analysis
Conjoint analysis
We test H-H using choice-based conjoint analysis (Hainmueller et al., 2014). Choice-based analysis is a technique that solicits user preferences over two options with various factors set to different attribute levels. The analysis generates the marginal means of each attribute level. Our analysis uses the cregg package in R (Leeper, 2020). For H, we look directly at the marginal mean effect of each of the multiplicity resolution techniques and costs. For each condition, the marginal mean procedure outputs a point estimate and a confidence interval of how preferred an option is relative to random chance. For instance, an output of 0.8 indicates that participants presented with that option chose it 80% of the time. For H, H, and H, we compute the marginal means stratified by task characteristics. To the best of our knowledge, there is no tool to compute marginal means for a continuous factor, so we use the binary encodings (low vs. high) for task stakes and task uncertainty.
Testing H
We will consider H to be weakly supported if the confidence interval for human is strictly greater than 0.5 and the confidence interval for random is strictly less than 0.5. Then, we will test whether the estimate for human is strictly higher than the estimates for the other methods. To do this, we will perform five two-sided t-tests555The choice-based analysis data may not be normally distributed; however, we have over 800 observations for each resolution technique so the central limit theorem makes the use of this test acceptable. (one against each other multiplicity resolution technique). Then, we will perform four t-tests (random vs. each other multiplicity resolution technique other than human) to see whether random is strictly less preferred than each other technique.
Testing H-H
We compute the marginal means model for each task characteristic. We then perform post-hoc t-tests on the estimates. For instance, to test H, we first check whether the estimate for human for low-stake tasks is different from its estimate for high-stake tasks. In total, we perform four post-hoc tests for H, three for H, and two for H.
Multiple hypothesis correction
Since there are six multiplicity resolution techniques, we use the Bonferroni666Note, we do not use Holm-Bonferroni here because it is not supported by cregg. correction to get a p-value of and use this to compute the confidence intervals for each marginal means estimate. We perform a total of 18 post-hoc t-tests on the data, so will evaluate their significance using the Holm-Bonferroni correction with and .
5. Data Analysis
| Description | Sig? | |
| H | Multiplicity reduces fairness perceptions | No |
| H | H is modulated by task stakes | No |
| H | H is modulated by task framing | No |
| H | H is modulated by task uncertainty | No |
| H | Participants care how multiplicity is resolved | Yes |
| H | H is modulated by task stakes | Yes |
| H | H is modulated by task framing | Yes |
| H | H is modulated by task uncertainty | No |
5.1. Summary of responses
We receive 367 survey responses and after excluding ones that fail the initial comprehension check twice or fail multiple attention checks, we use 358 in our analysis. 58.7% of respondents are women, 2.0% are non-binary, and 39.4% are men. The average respondent age is 36.3 (range: 18-78, stdev: 12.3). For the “pre” timestamp, Cronbach’s Alpha between the four Likert items is 0.88 and for the “post” timestamp it is 0.94.
5.2. Evaluation of hypotheses
We find no evidence for H (. H (), H (, or H ().
H is significant since different multiplicity resolution techniques are selected at statistically significantly different rates (see fig. 3). For instance, human is selected in a pairwise comparison at a rate of 0.795 (SE = 0.015, p ¡ 0.0001) and random is selected at a rate of 0.279 (SE = 0.017, p ¡ 0.0001). A post-hoc test (detailed results in appendix) reveals that the preference for human is statistically stronger than all other options, while random is significantly less preferred than all options other than default-bad.
H is significant as well, and one of our four specific predictions holds: human is more strongly preferred when a task is high stakes. H is significant, as is one of our four predictions: when the task is reward-based, default-good is more strongly preferred than when the task is punishment-based. H is not significant. Marginal mean estimates and details on t-test results are in the appendix.
5.3. Exploratory analyses
We describe five exploratory results (labeled E1-E5).
E1 (Differences between different dimensions of fairness)
While the set of 4 Likert items has high internal agreement for both the “pre” and “post” timestamps, the changes that occur with time are not distributed randomly across the four items. Overall, there is not a significant difference between “pre” and “post” fairness scores (H). But the first two Likert items, which ask about model outcomes (whether a system will be able to make good/fair decisions), see the scores decrease by an average of 0.122 points per item, while the second two Likert items, which ask about the appropriateness of using the AI system, increase with time by an average of 0.131 points per item. This difference is statistically significant according to the Mann-Whitney U-test (W=74118, p=0.0001).
The plot shows the marginal mean estimates for each multiplicity-resolution technique. We will report the lower bound, mean estimate, and upper bound for each. Default bad: 0.276, 0.328, 0.380. Default good: 0.522, 0.576, 0.630. Fancy: 0.551, 0.602, 0.653. Ignore: 0.367, 0.41, 0.464. Human: 0.749, 0.790, 0.832. Random: 0.235, 0.279, 0.323.
E2 (Survey duration’s impact on change in fairness scores)
We compute Spearman’s rank correlation between survey duration and average change in fairness score. There a very weak positive correlation (, p=0.0002).
E3 (Trends in multiplicity solution preferences)
We observe several trends that our hypotheses did not predict. (All results in the following paragraph are for t-tests based on marginal means estimates.) When stratifying tasks by stakes, default-bad is more preferred in low-stakes settings than in high-stakes settings (difference in marginal means -0.166, SE=0.0386, t=-4.307, p¡). We hypothesized that random would be more preferred in low-stakes solutions (H); this was not significant with the multiple hypothesis correction but shows a trend in the right direction (difference -0.079, SE=0.033, t=-2.411, p=0.016). The default-bad option is also more preferred in punishment settings than in reward settings (difference -0.183, SE=0.039, t=-4.692, p¡), while random is more preferred in reward settings (difference 0.108, SE=0.033, t=3.31, p=0.0009). In high-uncertainty settings, default-good and default-bad are both more preferred than in low-uncertainty settings (difference 0.108, SE=0.044, t=2.441, p=0.015 and difference 0.089, SE=0.0387, t=2.310, p=0.021, respectively).
E4 (Role of cost)
A marginal-means analysis shows that no extra cost is usually significantly preferred over extra cost (t-test result: difference 0.225, SE=0.016, t=13.91, p). However, in three settings (tumor detection, care plan access, and job probation), the marginal means for no extra cost and extra cost are not statistically different (see the appendix for complete data).
E5 (Qualitative analysis of open-ended question)
We analyze the responses to the open-ended question as follows: First, we exclude responses that are unclear or do not suggest a specific decision-making strategy. Then, we categorize responses according to the resolution techniques from table 1 and an “other” category. We then review the responses in the “other” category and group similar responses into new categories. If a response mentions multiple approaches, we count it in each category. Note that at this stage of the survey, participants had not seen the options in table 1. We report general trends here; full statistics are available in the appendix.
The majority of participants across all tasks mention human, including over 80% of participants for job probation, care plan access, and tumor detection. default-good is mentioned most frequently for tumor detection, tax audit, and beta app access, while default-bad is suggested for beta app access and flight upgrade. random is suggested frequently for flight upgrade and occasionally for tax audit and flight rebooking. Participants also frequently suggested specific rules they would implement or indicated that a more complex model could be built to resolve disagreements. This type of answer was especially common for flight upgrades and rebooking, review moderation, and beta app access. Finally, a few participants suggested involving the decision subject directly, e.g., asking a flight passenger how they feel about being rebooked or upgraded. This option was mentioned by at least one respondent for flight rebooking, flight upgrade, and care plan access.
6. Discussion
6.1. The standard ML approach is unfair (ignore is disliked)
Our most impactful finding is that ignore is generally not preferred.
Implication: Standard ML workflows need to adapt to consider multiplicity, such as by increasing transparency when multiplicity impacts a decision.
Depending on the target audience, interventions to increase transparency could take different forms. For instance, common machine-learning packages could alert developers when multiplicity exists (e.g., by automatically performing a bootstrapping analysis in the background). Model documentation should include an analysis of what proportion of test samples are subject to receive a conflicting prediction under multiplicity, so that anyone who decides to use or adapt the model in the future is aware. In public communications about a model, the fraction of inputs (based on validation data) that are subject to receive a different prediction under multiplicity could be reported.
Implication: Model developers should be able to justify how multiplicity is handled, whether this is by deferring to a human decision-maker, employing randomization, using ensemble methods, or something else.
It is possible that decision subjects would be okay with using a single model, and knowingly ignoring multiplicity, if this decision is justified. But, just as model developers need to justify decisions such as feature selection and statistical fairness trade-offs, they should justify why the existence of multiplicity does not pose an existential threat to their project. In particular, they should consider potential harms to individuals, e.g., whether the model’s use will be widespread and lead to systemic exclusion of some people.
6.2. Is multiplicity an inherent fairness risk?
Our results show that participants’ fairness perceptions did not decrease when multiplicity was present (i.e., H was not supported). This contradicts work in ML and philosophy that assumes and argues that multiplicity poses a fairness risk (see section 2.2). Since H was not supported by our data, either (1) the experts are incorrect about the fairness risks of multiplicity, (2) lay stakeholders do not see fairness risks of multiplicity due to low technical literacy, or (3) our study design was insufficient to capture stakeholders’ true views on multiplicity. We discuss (3) when describing our work’s limitations (section 6.6), so we will overview (1) and (2) here.
If (1), the fairness risks of multiplicity summarized in section 2.2 must paint an incomplete picture. Perhaps the philosophical view that only systematic arbitrariness is a fairness risk (Creel and Hellman, 2022) is correct. Alternatively, maybe multiplicity poses a theoretical fairness risk, but one that is is either too small to matter in practice, or that is overpowered by the benefits of using ML. Participants may feel as if the ML system is a closed box, and thus that its decisions are impenetrable. This perception would, understandably, lead participants to feel disenfranchised in their interactions with the ML model. From this perspective, multiplicity is not the primary reason why a model’s usage could be unfair.
Implication: Theoretical researchers should articulate how the fairness risks of multiplicity relate to other fairness issues in ML – such as accountability and recourse – to help clarify when, exactly, multiplicity poses a fairness risk. Future research about the perceived fairness of multiplicity should do more to access the “why:” what are the driving forces in fairness perceptions?
If (2), then H was not supported because the stakeholders’ technical literacy was not high enough to realize that multiplicity poses the procedural fairness risks suggested by computer scientists and philosophers. As ML technology powers more and more decisions that used to be made by humans, it is important that stakeholders are able to recognize its limitations to avoid inappropriate reliance on and trust in ML systems (Long and Magerko, 2020). One way that reliance and trust can be calibrated is through improving technical/AI literacy (Chiang and Yin, 2022). We discuss how our survey design could be modified to promote better understanding in section 6.6. But also, technical literacy can be built gradually over time by exposure to news stories and public discourse (Nguyen and Hekman, 2024).
Implication: We encourage people working at the intersection of technology and communication to discuss multiplicity and to take a critical lens, more broadly, when describing new ML systems in socially-impactful domains.
6.3. Randomization is disliked
Scholars in ML and philosophy typically suggest using a more complex model (e.g., via ensembles) (Cooper et al., 2024; Roth et al., 2023) or randomization (Creel and Hellman, 2022; Jain et al., 2024). Participants in the survey expressed comfort with fancy – choosing this option more often than would be expected by random chance – but strongly disliked random (H). This contradicts the established philosophical view that lotteries are fair in situations where a scarce resource must be allocated (Broome, 1990; Sher, 1980). A possible discrepancy between our setup and a run-of-the-mill lottery is the inclusion of machine learning in the decision-making process. Perhaps randomization via a lottery is familiar, but the addition of machine learning suggests that there must be some way to “learn” the best allocation, thereby making randomization too crude of a tool.
Implication: More research should be conducted into why random is disliked, even though lotteries are generally viewed as fair. Researchers can investigate whether providing additional education about machine learning (e.g., how models can make mistakes and that there is often not a single, optimal solution) improves comfort with random.
6.4. Participants prefer human intervention, but is that scalable?
Participants generally prefer human over other multiplicity-resolution techniques (H). However, when large proportions of predictions can vary under multiplicity, deferring to a human decision-maker is difficult to operationalize and negates cost and time savings associated with using ML. In these cases, one option is to simply forgo using ML at all to avoid the appearance of automation and the expense of developing a system that only handles a small number of cases. (Indeed, it may not be scientifically valid to use machine learning when many predictions vary under multiplicity, as has been argued for “predictive optimization” tasks (Wang et al., 2024).) Instead, decision-makers could rely on an interpretable, publicly-available set of guidelines. Alternatively, model developers could use a strategy that combines fancy and human. Recall that fancy (e.g., using ensemble methods) is favored by ML researchers as a technique to reduce multiplicity. Model developers could use ensembles or more complex model architectures to reduce multiplicity, and then defer to a human decision maker only for the smaller number of cases whose predictions remain ambiguous. In our experiments, participants were generally comfortable with fancy, so while further research is needed to confirm that a combination fancy and human approach would be viewed as fair, we are optimistic about that possibility.
Implication: Model developers may not be able to align with stakeholders’ preferences for human in all cases, but should consider whether other, more interpretable decision-making rules are more appropriate when multiplicity is present. Alternatively, using ensemble methods to reduce multiplicity, and then deferring to a human on the smaller number of ambiguous cases, could strike a good balance.
6.5. Impact of task choice
Preferences for multiplicity resolution techniques varied across tasks. E3 shows that random is more preferred in general for low-stakes scenarios, but also for one high-stakes scenario (tax audit). We hypothesize that this is because decisions about whom to audit are seen as more impersonal than the other high-stakes scenarios. In medical and employment domains, decision subjects typically have a personal relationship with their doctor or manager, and thus trust that they will have the ability to appeal to human emotion and get an explanation of the decision. By contrast, decisions by the IRS (the US tax agency) about whom to audit likely already feel random to many individuals. Conversely, human is more strongly preferred in high-stakes settings where a personal connection is the norm (H & E3). We also want to call attention to how ignore is usually preferred to random (H). From the perspective of a decision subject who interacts with a model a single time, there is not a mathematical difference between ignore and random because the personal outcome is unknown and (relatively) unintelligent in each. (By contrast, default-good and default-bad give a known outcome, and people may assume that they can appeal to a human decision maker to get the correct/favorable outcome with human and that fancy will be smart enough to give the correct decision.) So, individuals must either be imagining repeated encounters with a system or just view having a (presumably) deterministic model as providing benefits such as accountability and predictability.
Implication: Model developers should involve participants in the design process to figure out how they prefer that multiplicity be handled for the specific task.
6.6. Limitations and opportunities for future work
We conclude our discussion with an overview of limitations and how they could be addressed in future work.
Cognitive load for participants
A limitation of this work is that multiplicity is not an easy concept to grasp, especially for people without machine learning experience. We attempted to explain it clearly and in simple terms and iterated on the survey text through multiple pilot iterations. Our results do not show significant differences in the responses for participants with or without ML experience. We attempted to check that participants understood multiplicity through an initial comprehension check, but some participants may have gotten lucky and passed this question by chance, only looked for the correct answer rather than engaging critically with the reading, or may have forgotten details when answering the remaining survey questions.
Unknown role of other task-related factors
To counteract the cognitive load required to reason about multiplicity, we tried to make the other details about the tasks and models as simple as possible, including by providing only high-level details of the AI system’s functionality and context. Although we tried to select tasks that are broadly relevant, some participants may not have seen the relevance to their own lives or imagined themselves as the decision subjects. Some participants may have felt like they did not have enough information to evaluate the system’s fairness, and the presence of multiplicity may have seemed insignificant relative to other unknown factors. In particular, we did not include any explanations of how the models made their decisions to keep the setup simpler. Informational transparency is a component of overall fairness (Colquitt, 2001), though, so future work could explore how different explanations of both model behavior and of multiplicity itself impact fairness perceptions. Likewise, we do not include personal impact – i.e., whether the prediction a model makes on a specific decision subject varies under multiplicity – as a factor due to the experiment’s complexity. We hypothesize – and future work should check – that if participants are told that the decision the model makes for them varies under multiplicity, fairness perceptions would decrease, mirroring work on outcome favorability (Wang et al., 2020). To combat issues with the large number of unknowns inherent to hypothetical situations, future research could be situated in a real-world task, where explaining more details about the model would not necessitate a large cognitive load due to baseline familiarity with an existing algorithm. Existing research also shows that people behave differently in hypothetical settings (FeldmanHall et al., 2012), again highlighting the need for additional research situated in a real use case.
Unknown role of other personal factors
We did not find significant trends between demographic groups for any of the factors we recorded. However, our analysis was not powerful enough to detect how any intersectional identities (e.g., gender and education level) may impact perceptions of multiplicity. Furthermore, our collection of some demographic attributes, especially experience with AI, may have been too coarse to see meaningful trends. Future work could consider contrasting specific participant groups (e.g., ML engineers versus adults with no ML experience) to gain a better sense of how fairness perceptions vary with ML experience.
Addressing the “why” behind multiplicity preferences
Our results indicate that stakeholders have strong preferences for how multiplicity is handled; however, our experiment does not determine why these preferences exist. In particular, the discrepancy between philosophical arguments for the fairness of random and the strong dislike that survey participants expressed towards random merits more attention. Is there truly a mismatch between what is theoretically fair based on philosophical principles and what people perceive as fair? Or is there a breakdown in the communication, e.g., would participants actually be comfortable with randomization if more details about the model and its operationalization were provided? Future studies could vary the amount of detail provided to see whether attitudes towards randomization change.
Task choice and characterization
We consider tasks that differ along three axes, but there may be alternate ways to characterize tasks. It is also not necessarily straightforward to introduce new tasks into our classification system, particularly for the continuous scale of task stakes and task uncertainty. Our analysis also lost detail about the task characteristics in the conjoint analysis, as this framework does not support continuous factors. Future work could continue to study how to systematically characterize tasks to promote reproducible and generalizable research.
US-centric perspective
We conducted our survey with US-based crowdworkers. It is known that fairness perceptions vary cross-culturally (Berman et al., 1985; Kim and Leung, 2007), as do attitudes towards uncertainty (Hofstede, 1984). It therefore seems likely that attitudes towards the fairness of multiplicity in ML would vary across cultures, too, which future work could explore.
7. Conclusions
We conducted a study (n=357) of crowdworkers to determine how lay stakeholders (1) perceive the fairness of ML models when multiplicity impacts decisions, and (2) prefer that this multiplicity be resolved. Our results indicate that stakeholders have strong preferences about how multiplicity should be resolved. In particular, they express a preference for human decision makers, and a dispreference for both ignoring multiplicity (arbitrarily using a single model) and randomizing between different models’ outputs. These findings counter not only the established ML practice of looking for a single well-performing model, but also philosophical arguments that randomization is the fairest way to deal with multiplicity. Despite these strong opinions about resolving multiplicity, the presence of multiplicity did not impact how fair the participants thought the models were, which is a discrepancy that future work will need to address. One promising avenue is to consider how perceptions of informational fairness relating to multiplicity, e.g., the use of explanations, intersects with the procedural components of fairness that we investigated. Ultimately, a greater understanding of when ML design choices – including how multiplicity is resolved – are perceived as fair can help guide model developers to create technology that aligns with stakeholders’ values.
Acknowledgements.
The authors thank members of the MadPL group for serving as pilot testers for initial survey iterations, and Michael Coblenz for his comments on an earlier draft. This work was conducted with the support of NSF grants 1918211 and 2446711.References
- (1)
- Araujo et al. (2020) Theo Araujo, Natali Helberger, Sanne Kruikemeier, and Claes H De Vreese. 2020. In AI we trust? Perceptions about automated decision-making by artificial intelligence. AI & society 35, 3 (2020), 611–623.
- Bankins et al. (2022) Sarah Bankins, Paul Formosa, Yannick Griep, and Deborah Richards. 2022. AI decision making with dignity? Contrasting workers’ justice perceptions of human and AI decision making in a human resource management context. Information Systems Frontiers 24, 3 (2022), 857–875.
- Bartlett and Wegkamp (2008) Peter L Bartlett and Marten H Wegkamp. 2008. Classification with a Reject Option using a Hinge Loss. Journal of Machine Learning Research 9, 8 (2008), 1823–1840.
- Bates et al. (2015) Douglas Bates, Martin Mächler, Ben Bolker, and Steve Walker. 2015. Fitting Linear Mixed-Effects Models Using lme4. Journal of Statistical Software 67, 1 (2015), 1–48. https://doi.org/10.18637/jss.v067.i01
- Berman et al. (1985) John J Berman, Virginia Murphy-Berman, and Purnima Singh. 1985. Cross-cultural similarities and differences in perceptions of fairness. Journal of Cross-Cultural Psychology 16, 1 (1985), 55–67.
- Binns et al. (2018) Reuben Binns, Max Van Kleek, Michael Veale, Ulrik Lyngs, Jun Zhao, and Nigel Shadbolt. 2018. ’It’s Reducing a Human Being to a Percentage’ Perceptions of Justice in Algorithmic Decisions. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3173574.3173951
- Black et al. (2022a) Emily Black, Klas Leino, and Matt Fredrikson. 2022a. Selective Ensembles for Consistent Predictions. In International Conference on Learning Representations. https://openreview.net/forum?id=HfUyCRBeQc
- Black et al. (2022b) Emily Black, Manish Raghavan, and Solon Barocas. 2022b. Model multiplicity: Opportunities, concerns, and solutions. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency. Association for Computing Machinery, New York, NY, USA, 850–863. https://doi.org/10.1145/3531146.3533149
- Blanca Mena et al. (2017) M José Blanca Mena, Rafael Alarcón Postigo, Jaume Arnau Gras, Roser Bono Cabré, and Rebecca Bendayan. 2017. Non-normal data: Is ANOVA still a valid option? Psicothema 29, 4 (2017), 552–557.
- Bondi et al. (2021) Elizabeth Bondi, Lily Xu, Diana Acosta-Navas, and Jackson A Killian. 2021. Envisioning communities: a participatory approach towards AI for social good. In Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society. Association for Computing Machinery, New York, NY, USA, 425–436. https://doi.org/10.1145/3461702.3462612
- Breiman (2001) Leo Breiman. 2001. Statistical modeling: The two cultures (with comments and a rejoinder by the author). Statistical science 16, 3 (2001), 199–231.
- Broome (1990) John Broome. 1990. Fairness. In Proceedings of the Aristotelian society, Vol. 91. JSTOR, [Aristotelian Society, Wiley], 87–101.
- Carifio and Perla (2008) James Carifio and Rocco Perla. 2008. Resolving the 50-year debate around using and misusing Likert scales. , 1150–1152 pages.
- Castelo et al. (2019) Noah Castelo, Maarten W Bos, and Donald R Lehmann. 2019. Task-dependent algorithm aversion. Journal of Marketing Research 56, 5 (2019), 809–825.
- Chiang and Yin (2022) Chun-Wei Chiang and Ming Yin. 2022. Exploring the Effects of Machine Learning Literacy Interventions on Laypeople’s Reliance on Machine Learning Models. In Proceedings of the 27th International Conference on Intelligent User Interfaces (Helsinki, Finland) (IUI ’22). Association for Computing Machinery, New York, NY, USA, 148–161. https://doi.org/10.1145/3490099.3511121
- Chouldechova (2017) Alexandra Chouldechova. 2017. Fair prediction with disparate impact: A study of bias in recidivism prediction instruments. Big Data 5, 2 (2017), 153–163.
- Chouldechova et al. (2018) Alexandra Chouldechova, Diana Benavides-Prado, Oleksandr Fialko, and Rhema Vaithianathan. 2018. A case study of algorithm-assisted decision making in child maltreatment hotline screening decisions. In Proceedings of the 1st Conference on Fairness, Accountability and Transparency (Proceedings of Machine Learning Research, Vol. 81), Sorelle A. Friedler and Christo Wilson (Eds.). PMLR, 134–148. https://proceedings.mlr.press/v81/chouldechova18a.html
- Cohen et al. (2019) Jeremy Cohen, Elan Rosenfeld, and Zico Kolter. 2019. Certified Adversarial Robustness via Randomized Smoothing. In Proceedings of the 36th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 97), Kamalika Chaudhuri and Ruslan Salakhutdinov (Eds.). PMLR, 1310–1320. https://proceedings.mlr.press/v97/cohen19c.html
- Colquitt (2001) Jason A Colquitt. 2001. On the dimensionality of organizational justice: a construct validation of a measure. Journal of applied psychology 86, 3 (2001), 386.
- Colquitt and Rodell (2015) Jason A Colquitt and Jessica B Rodell. 2015. Measuring justice and fairness. The Oxford Handbook of Justice in the Workplace 1 (2015), 187–202.
- Cooper et al. (2024) A. Feder Cooper, Katherine Lee, Madiha Zahrah Choksi, Solon Barocas, Christopher De Sa, James Grimmelmann, Jon Kleinberg, Siddhartha Sen, and Baobao Zhang. 2024. Arbitrariness and Social Prediction: The Confounding Role of Variance in Fair Classification. Proceedings of the AAAI Conference on Artificial Intelligence 38, 20 (Mar. 2024), 22004–22012. https://doi.org/10.1609/aaai.v38i20.30203
- Creel and Hellman (2022) Kathleen Creel and Deborah Hellman. 2022. The Algorithmic Leviathan: Arbitrariness, Fairness, and Opportunity in Algorithmic Decision-Making Systems. Canadian Journal of Philosophy 52, 1 (2022), 26–43. https://doi.org/10.1017/can.2022.3
- Cushman et al. (2006) Fiery Cushman, Liane Young, and Marc Hauser. 2006. The role of conscious reasoning and intuition in moral judgment: Testing three principles of harm. Psychological Science 17, 12 (2006), 1082–1089.
- D’Amour et al. (2022) Alexander D’Amour, Katherine Heller, Dan Moldovan, Ben Adlam, Babak Alipanahi, Alex Beutel, Christina Chen, Jonathan Deaton, Jacob Eisenstein, Matthew D Hoffman, et al. 2022. Underspecification Presents Challenges for Credibility in Modern Machine Learning. Journal of Machine Learning Research 23, 226 (2022), 1–61.
- Dietvorst and Bharti (2020) Berkeley J Dietvorst and Soaham Bharti. 2020. People reject algorithms in uncertain decision domains because they have diminishing sensitivity to forecasting error. Psychological Science 31, 10 (2020), 1302–1314.
- Dietvorst et al. (2015) Berkeley J Dietvorst, Joseph P Simmons, and Cade Massey. 2015. Algorithm aversion: people erroneously avoid algorithms after seeing them err. Journal of experimental psychology: General 144, 1 (2015), 114.
- Dodge et al. (2019) Jonathan Dodge, Q Vera Liao, Yunfeng Zhang, Rachel KE Bellamy, and Casey Dugan. 2019. Explaining models: an empirical study of how explanations impact fairness judgment. In Proceedings of the 24th International Conference on Intelligent User Interfaces. Association for Computing Machinery, New York, NY, USA, 275–285. https://doi.org/10.1145/3301275.3302310
- Dwork et al. (2012) Cynthia Dwork, Moritz Hardt, Toniann Pitassi, Omer Reingold, and Richard Zemel. 2012. Fairness through Awareness. In Proceedings of the 3rd Innovations in Theoretical Computer Science Conference (Cambridge, Massachusetts) (ITCS ’12). Association for Computing Machinery, 214–226. https://doi.org/10.1145/2090236.2090255
- Elkin et al. (2021) Lisa A Elkin, Matthew Kay, James J Higgins, and Jacob O Wobbrock. 2021. An aligned rank transform procedure for multifactor contrast tests. In The 34th annual ACM symposium on user interface software and technology. 754–768.
- Feffer et al. (2023) Michael Feffer, Michael Skirpan, Zachary Lipton, and Hoda Heidari. 2023. From preference elicitation to participatory ML: A critical survey & guidelines for future research. In Proceedings of the 2023 AAAI/ACM Conference on AI, Ethics, and Society. Association for Computing Machinery, New York, NY, USA, 38–48. https://doi.org/10.1145/3600211.3604661
- FeldmanHall et al. (2012) Oriel FeldmanHall, Dean Mobbs, Davy Evans, Lucy Hiscox, Lauren Navrady, and Tim Dalgleish. 2012. What we say and what we do: The relationship between real and hypothetical moral choices. Cognition 123, 3 (2012), 434–441.
- Fogliato et al. (2022) Riccardo Fogliato, Shreya Chappidi, Matthew Lungren, Paul Fisher, Diane Wilson, Michael Fitzke, Mark Parkinson, Eric Horvitz, Kori Inkpen, and Besmira Nushi. 2022. Who goes first? Influences of human-AI workflow on decision making in clinical imaging. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency. Association for Computing Machinery, New York, NY, USA, 1362–1374. https://doi.org/10.1145/3531146.3533193
- Fox and Weisberg (2019) John Fox and Sanford Weisberg. 2019. An R Companion to Applied Regression (third ed.). Sage, Thousand Oaks CA. https://socialsciences.mcmaster.ca/jfox/Books/Companion/
- Grgic-Hlaca et al. (2018) Nina Grgic-Hlaca, Elissa M. Redmiles, Krishna P. Gummadi, and Adrian Weller. 2018. Human Perceptions of Fairness in Algorithmic Decision Making: A Case Study of Criminal Risk Prediction. In Proceedings of the 2018 World Wide Web Conference (Lyon, France) (WWW ’18). International World Wide Web Conferences Steering Committee, Republic and Canton of Geneva, CHE, 903–912. https://doi.org/10.1145/3178876.3186138
- Hainmueller et al. (2014) Jens Hainmueller, Daniel J Hopkins, and Teppei Yamamoto. 2014. Causal inference in conjoint analysis: Understanding multidimensional choices via stated preference experiments. Political analysis 22, 1 (2014), 1–30.
- Hannan et al. (2021) Jacqueline Hannan, Huei-Yen Winnie Chen, and Kenneth Joseph. 2021. Who gets what, according to whom? An analysis of fairness perceptions in service allocation. In Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society. Association for Computing Machinery, New York, NY, USA, 555–565. https://doi.org/10.1145/3461702.3462568
- Hardt et al. (2016) Moritz Hardt, Eric Price, and Nati Srebro. 2016. Equality of opportunity in supervised learning. Advances in Neural Information Processing Systems 29 (2016).
- Helberger et al. (2020) Natali Helberger, Theo Araujo, and Claes H de Vreese. 2020. Who is the fairest of them all? Public attitudes and expectations regarding automated decision-making. Computer Law & Security Review 39 (2020), 105456.
- Herbei and Wegkamp (2006) Radu Herbei and Marten H Wegkamp. 2006. Classification with reject option. The Canadian Journal of Statistics/La Revue Canadienne de Statistique 34, 4 (2006), 709–721.
- Hofstede (1984) Geert Hofstede. 1984. Culture’s consequences: International differences in work-related values. Vol. 5. sage.
- Holm (1979) Sture Holm. 1979. A simple sequentially rejective multiple test procedure. Scandinavian journal of statistics 6, 2 (1979), 65–70.
- Howard et al. (1979) George S Howard, Kenneth M Ralph, Nancy A Gulanick, Scott E Maxwell, Don W Nance, and Sterling K Gerber. 1979. Internal invalidity in pretest-posttest self-report evaluations and a re-evaluation of retrospective pretests. Applied Psychological Measurement 3, 1 (1979), 1–23.
- Jain et al. (2024) Shomik Jain, Kathleen Creel, and Ashia Camage Wilson. 2024. Position: Scarce Resource Allocations That Rely On Machine Learning Should Be Randomized. In Forty-first International Conference on Machine Learning. https://openreview.net/forum?id=44qxX6Ty6F
- Jakesch et al. (2022) Maurice Jakesch, Zana Buçinca, Saleema Amershi, and Alexandra Olteanu. 2022. How different groups prioritize ethical values for responsible AI. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency. Association for Computing Machinery, New York, NY, USA, 310–323. https://doi.org/10.1145/3531146.3533097
- Jamieson (2004) Susan Jamieson. 2004. Likert scales: How to (ab) use them? Medical education 38, 12 (2004), 1217–1218.
- Kamiran and Calders (2012) Faisal Kamiran and Toon Calders. 2012. Data preprocessing techniques for classification without discrimination. Knowledge and Information Systems 33, 1 (2012), 1–33.
- Kasinidou et al. (2021) Maria Kasinidou, Styliani Kleanthous, Pınar Barlas, and Jahna Otterbacher. 2021. I agree with the decision, but they didn’t deserve this: Future Developers’ Perception of Fairness in Algorithmic Decisions. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency (Virtual Event, Canada) (FAccT ’21). Association for Computing Machinery, New York, NY, USA, 690–700. https://doi.org/10.1145/3442188.3445931
- Kay et al. (2021) Matthew Kay, Lisa A. Elkin, James J. Higgins, and Jacob O. Wobbrock. 2021. mjskay/ARTool: ARTool 0.11.0. https://doi.org/10.5281/zenodo.4721941
- Kim and Leung (2007) Tae-Yeol Kim and Kwok Leung. 2007. Forming and reacting to overall fairness: A cross-cultural comparison. Organizational Behavior and Human Decision Processes 104, 1 (2007), 83–95.
- Kleinberg et al. (2017) Jon Kleinberg, Sendhil Mullainathan, and Manish Raghavan. 2017. Inherent Trade-Offs in the Fair Determination of Risk Scores. In 8th Innovations in Theoretical Computer Science Conference (ITCS 2017). Schloss-Dagstuhl-Leibniz Zentrum für Informatik, Schloss Dagstuhl – Leibniz-Zentrum für Informatik, Dagstuhl, Germany. https://drops.dagstuhl.de/entities/document/10.4230/LIPIcs.ITCS.2017.43
- Kleinberg and Raghavan (2021) Jon Kleinberg and Manish Raghavan. 2021. Algorithmic monoculture and social welfare. Proceedings of the National Academy of Sciences 118, 22 (2021), e2018340118.
- Kusner et al. (2017) Matt J Kusner, Joshua Loftus, Chris Russell, and Ricardo Silva. 2017. Counterfactual fairness. Advances in Neural Information Processing Systems 30 (2017).
- Kuzon et al. (1996) William Kuzon, Melanie Urbanchek, and Steven McCabe. 1996. The seven deadly sins of statistical analysis. Annals of plastic surgery 37 (1996), 265–272.
- Laidlaw and Feizi (2019) Cassidy Laidlaw and Soheil Feizi. 2019. Playing it safe: Adversarial robustness with an abstain option. arXiv preprint arXiv:1911.11253 (2019).
- Lee (2018) Min Kyung Lee. 2018. Understanding perception of algorithmic decisions: Fairness, trust, and emotion in response to algorithmic management. Big Data & Society 5, 1 (2018), 2053951718756684.
- Lee and Rich (2021) Min Kyung Lee and Katherine Rich. 2021. Who Is Included in Human Perceptions of AI?: Trust and Perceived Fairness around Healthcare AI and Cultural Mistrust. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 138, 14 pages. https://doi.org/10.1145/3411764.3445570
- Lee et al. (2024) Yoonjoo Lee, Kihoon Son, Tae Soo Kim, Jisu Kim, John Joon Young Chung, Eytan Adar, and Juho Kim. 2024. One vs. Many: Comprehending Accurate Information from Multiple Erroneous and Inconsistent AI Generations. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. Association for Computing Machinery, New York, NY, USA, 2518–2531. https://doi.org/10.1145/3630106.3662681
- Leeper (2020) Thomas J. Leeper. 2020. cregg: Simple Conjoint Analyses and Visualization. R package version 0.4.0.
- Li et al. (2021) Lan Li, Tina Lassiter, Joohee Oh, and Min Kyung Lee. 2021. Algorithmic hiring in practice: Recruiter and HR Professional’s perspectives on AI use in hiring. In Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society. Association for Computing Machinery, New York, NY, USA, 166–176. https://doi.org/10.1145/3461702.3462531
- Logg et al. (2019) Jennifer M Logg, Julia A Minson, and Don A Moore. 2019. Algorithm appreciation: People prefer algorithmic to human judgment. Organizational Behavior and Human Decision Processes 151 (2019), 90–103.
- Long and Magerko (2020) Duri Long and Brian Magerko. 2020. What is AI Literacy? Competencies and Design Considerations. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3313831.3376727
- Luce (1977) R.Duncan Luce. 1977. The choice axiom after twenty years. Journal of Mathematical Psychology 15, 3 (1977), 215–233. https://doi.org/10.1016/0022-2496(77)90032-3
- Mann and Whitney (1947) Henry B Mann and Donald R Whitney. 1947. On a test of whether one of two random variables is stochastically larger than the other. The Annals of Mathematical Statistics 18, 1 (1947), 50–60.
- Marx et al. (2020) Charles Marx, Flavio Calmon, and Berk Ustun. 2020. Predictive Multiplicity in Classification. In Proceedings of the 37th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 119), Hal Daumé III and Aarti Singh (Eds.). PMLR, 6765–6774. https://proceedings.mlr.press/v119/marx20a.html
- Maystre and Grossglauser (2015) Lucas Maystre and Matthias Grossglauser. 2015. Fast and accurate inference of Plackett–Luce models. Advances in Neural Information Processing Systems 28 (2015).
- Menon and Williamson (2018) Aditya Krishna Menon and Robert C Williamson. 2018. The cost of fairness in binary classification. In Conference on Fairness, accountability and transparency (Proceedings of Machine Learning Research, Vol. 81), Sorelle A. Friedler and Christo Wilson (Eds.). PMLR, PMLR, 107–118. https://proceedings.mlr.press/v81/menon18a.html
- Narayanan et al. (2024) Devesh Narayanan, Mahak Nagpal, Jack McGuire, Shane Schweitzer, and David De Cremer. 2024. Fairness perceptions of artificial intelligence: A review and path forward. International Journal of Human–Computer Interaction 40, 1 (2024), 4–23.
- Nguyen and Hekman (2024) Dennis Nguyen and Erik Hekman. 2024. The news framing of artificial intelligence: a critical exploration of how media discourses make sense of automation. AI & Society 39, 2 (2024), 437–451.
- Norman (2010) Geoff Norman. 2010. Likert scales, levels of measurement and the “laws” of statistics. Advances in health sciences education 15 (2010), 625–632.
- Obermeyer et al. (2019) Ziad Obermeyer, Brian Powers, Christine Vogeli, and Sendhil Mullainathan. 2019. Dissecting racial bias in an algorithm used to manage the health of populations. Science 366, 6464 (2019), 447–453.
- Plackett (1975) R. L. Plackett. 1975. The Analysis of Permutations. Journal of the Royal Statistical Society. Series C (Applied Statistics) 24, 2 (1975), 193–202. http://www.jstor.org/stable/2346567
- Prolific (2024) Prolific. 2024. Prolific’s Attention and Comprehension Check Policy. https://researcher-help.prolific.com/en/article/fb63bb
- Propublica (2016) Propublica. 2016. Propublica Compas Analysis. https://github.com/propublica/compas-analysis
- Roth et al. (2023) Aaron Roth, Alexander Tolbert, and Scott Weinstein. 2023. Reconciling Individual Probability Forecasts*. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (¡conf-loc¿, ¡city¿Chicago¡/city¿, ¡state¿IL¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 101–110. https://doi.org/10.1145/3593013.3593980
- Rudin et al. (2024) Cynthia Rudin, Chudi Zhong, Lesia Semenova, Margo Seltzer, Ronald Parr, Jiachang Liu, Srikar Katta, Jon Donnelly, Harry Chen, and Zachery Boner. 2024. Position: Amazing Things Come From Having Many Good Models. In Forty-first International Conference on Machine Learning. https://openreview.net/forum?id=oFDFGd9Age
- Sagi and Rokach (2018) Omer Sagi and Lior Rokach. 2018. Ensemble learning: A survey. Wiley interdisciplinary reviews: data mining and knowledge discovery 8, 4 (2018), e1249.
- Salimzadeh et al. (2023) Sara Salimzadeh, Gaole He, and Ujwal Gadiraju. 2023. A missing piece in the puzzle: Considering the role of task complexity in human-ai decision making. In Proceedings of the 31st ACM Conference on User Modeling, Adaptation and Personalization. Association for Computing Machinery, New York, NY, USA, 215–227. https://doi.org/10.1145/3565472.3592959
- Salimzadeh et al. (2024) Sara Salimzadeh, Gaole He, and Ujwal Gadiraju. 2024. Dealing with Uncertainty: Understanding the Impact of Prognostic Versus Diagnostic Tasks on Trust and Reliance in Human-AI Decision-Making. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 17 pages. https://doi.org/10.1145/3613904.3641905
- Saragih and Morrison (2022) Melissa Saragih and Ben W Morrison. 2022. The effect of past algorithmic performance and decision significance on algorithmic advice acceptance. International Journal of Human–Computer Interaction 38, 13 (2022), 1228–1237.
- Saxena et al. (2019) Nripsuta Ani Saxena, Karen Huang, Evan DeFilippis, Goran Radanovic, David C. Parkes, and Yang Liu. 2019. How Do Fairness Definitions Fare? Examining Public Attitudes Towards Algorithmic Definitions of Fairness. In Proceedings of the 2019 AAAI/ACM Conference on AI, Ethics, and Society (Honolulu, HI, USA) (AIES ’19). Association for Computing Machinery, New York, NY, USA, 99–106. https://doi.org/10.1145/3306618.3314248
- Schoeffer et al. (2022) Jakob Schoeffer, Niklas Kuehl, and Yvette Machowski. 2022. “There is not enough information”: On the effects of explanations on perceptions of informational fairness and trustworthiness in automated decision-making. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency. Association for Computing Machinery, New York, NY, USA, 1616–1628. https://doi.org/10.1145/3531146.3533218
- Scott et al. (2021) Ian A Scott, Stacy M Carter, and Enrico Coiera. 2021. Exploring stakeholder attitudes towards AI in clinical practice. BMJ Health & Care Informatics 28, 1 (2021).
- Searle et al. (1980) Shayle R Searle, F Michael Speed, and George A Milliken. 1980. Population marginal means in the linear model: an alternative to least squares means. The American Statistician 34, 4 (1980), 216–221.
- Sela (2018) Ayelet Sela. 2018. Can computers be fair: How automated and human-powered online dispute resolution affect procedural justice in mediation and arbitration. Ohio St. J. on Disp. Resol. 33 (2018), 91.
- Selbst et al. (2019) Andrew D Selbst, Danah Boyd, Sorelle A Friedler, Suresh Venkatasubramanian, and Janet Vertesi. 2019. Fairness and abstraction in sociotechnical systems. In Proceedings of the Conference on Fairness, Accountability, and Transparency. Association for Computing Machinery, New York, NY, USA, 59–68. https://doi.org/10.1145/3287560.3287598
- Sher (1980) George Sher. 1980. What makes a lottery fair? Noûs 14, 2 (1980), 203–216.
- Shin and Park (2019) Donghee Shin and Yong Jin Park. 2019. Role of fairness, accountability, and transparency in algorithmic affordance. Computers in Human Behavior 98 (2019), 277–284.
- Shin et al. (2020) Donghee Shin, Bu Zhong, and Frank A Biocca. 2020. Beyond user experience: What constitutes algorithmic experiences? International Journal of Information Management 52 (2020), 102061.
- Simmons et al. (2011) Joseph P Simmons, Leif D Nelson, and Uri Simonsohn. 2011. False-positive psychology: Undisclosed flexibility in data collection and analysis allows presenting anything as significant. Psychological science 22, 11 (2011), 1359–1366.
- Srivastava et al. (2019) Megha Srivastava, Hoda Heidari, and Andreas Krause. 2019. Mathematical Notions vs. Human Perception of Fairness: A Descriptive Approach to Fairness for Machine Learning. In Proceedings of the 25th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining (Anchorage, AK, USA) (KDD ’19). Association for Computing Machinery, New York, NY, USA, 2459–2468. https://doi.org/10.1145/3292500.3330664
- Starke et al. (2022) Christopher Starke, Janine Baleis, Birte Keller, and Frank Marcinkowski. 2022. Fairness perceptions of algorithmic decision-making: A systematic review of the empirical literature. Big Data & Society 9, 2 (2022), 20539517221115189.
- Steegen et al. (2016) Sara Steegen, Francis Tuerlinckx, Andrew Gelman, and Wolf Vanpaemel. 2016. Increasing transparency through a multiverse analysis. Perspectives on Psychological Science 11, 5 (2016), 702–712.
- Umbrello and Van de Poel (2021) Steven Umbrello and Ibo Van de Poel. 2021. Mapping value sensitive design onto AI for social good principles. AI and Ethics 1, 3 (2021), 283–296.
- Wang et al. (2024) Angelina Wang, Sayash Kapoor, Solon Barocas, and Arvind Narayanan. 2024. Against predictive optimization: On the legitimacy of decision-making algorithms that optimize predictive accuracy. ACM Journal on Responsible Computing 1, 1 (2024), 1–45.
- Wang et al. (2020) Ruotong Wang, F Maxwell Harper, and Haiyi Zhu. 2020. Factors influencing perceived fairness in algorithmic decision-making: Algorithm outcomes, development procedures, and individual differences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3313831.3376813
- Watson-Daniels et al. (2023) Jamelle Watson-Daniels, David C. Parkes, and Berk Ustun. 2023. Predictive Multiplicity in Probabilistic Classification. Proceedings of the AAAI Conference on Artificial Intelligence 37, 9 (Jun. 2023), 10306–10314. https://doi.org/10.1609/aaai.v37i9.26227
- Wobbrock et al. (2011) Jacob O Wobbrock, Leah Findlater, Darren Gergle, and James J Higgins. 2011. The aligned rank transform for nonparametric factorial analyses using only anova procedures. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 143–146. https://doi.org/10.1145/1978942.1978963
- Woodruff et al. (2018) Allison Woodruff, Sarah E Fox, Steven Rousso-Schindler, and Jeffrey Warshaw. 2018. A qualitative exploration of perceptions of algorithmic fairness. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3173574.3174230
- Wu et al. (2022) Xingjiao Wu, Luwei Xiao, Yixuan Sun, Junhang Zhang, Tianlong Ma, and Liang He. 2022. A survey of human-in-the-loop for machine learning. Future Generation Computer Systems 135 (2022), 364–381.
- Xin et al. (2022) Rui Xin, Chudi Zhong, Zhi Chen, Takuya Takagi, Margo Seltzer, and Cynthia Rudin. 2022. Exploring the Whole Rashomon Set of Sparse Decision Trees. In Advances in Neural Information Processing Systems, Alice H. Oh, Alekh Agarwal, Danielle Belgrave, and Kyunghyun Cho (Eds.). Curran Associates, Inc., 14071–14084. https://openreview.net/forum?id=WHqVVk3UHr
- Yin et al. (2019) Ming Yin, Jennifer Wortman Vaughan, and Hanna Wallach. 2019. Understanding the effect of accuracy on trust in machine learning models. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3290605.3300509
- Yurrita et al. (2023) Mireia Yurrita, Tim Draws, Agathe Balayn, Dave Murray-Rust, Nava Tintarev, and Alessandro Bozzon. 2023. Disentangling Fairness Perceptions in Algorithmic Decision-Making: the Effects of Explanations, Human Oversight, and Contestability. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 134, 21 pages. https://doi.org/10.1145/3544548.3581161
- Zhang et al. (2023) Angie Zhang, Olympia Walker, Kaci Nguyen, Jiajun Dai, Anqing Chen, and Min Kyung Lee. 2023. Deliberating with AI: Improving Decision-Making for the Future through Participatory AI Design and Stakeholder Deliberation. Proc. ACM Hum.-Comput. Interact. 7, CSCW1, Article 125 (apr 2023), 32 pages. https://doi.org/10.1145/3579601
Appendix A Preliminary study
A.1. Survey materials
We ask the following questions (on 5-point Likert scales from strongly disagree to strongly agree) to gauge task stakes (S), task uncertainty (C), and task clarity and realistic-ness (R). Questions are flagged with (*) to indicate that the phrasing changes with the scenario to provide more detail (e.g., for the task of loan approval “in the positive direction” would become “approves a loan for someone who will not be able to repay it”, since getting a loan is the positive outcome from the perspective of the decision subject).
(S) This scenario describes a high-stakes situation.
(S) It is a big problem if the system makes a mistake in the positive direction.*
(S) It is a big problem if the system makes a mistake in the negative direction.*
(C) For each decision the system makes, it is objectively right or wrong (i.e., there is not much ambiguity in decisions).
(C) If the AI system is designed well, I think it would be able to make better decisions than a human expert.
(R) Is this scenario described clearly?
(R) Is this scenario realistic?
A.2. Tasks and results
Table 4 shows all tasks we considered in the preliminary study, along with their final stakes and uncertainty scores.
| Domain | Task | Stakes | Uncertainty | Framing | Use |
| Healthcare | Access to a preventative health care plan | 0.42 | 0.91 | Reward | Yes |
| Business | Job probation based on employee monitoring | 0.72 | 1.12 | Punishment | Yes |
| Healthcare | Tumor detection from medical imaging | 1.61 | -1.37 | Reward | Yes |
| Government | Tax audit | 0.44 | -1.34 | Punishment | Yes |
| Travel | Airline seat upgrade | -1.94 | 1.18 | Reward | Yes |
| Travel | Rebooking passengers on an over-full flight | -0.27 | 0.50 | Punishment | Yes |
| Business | Customers get beta access to an app with discounts | -1.58 | -0.21 | Reward | Yes |
| Business | Removing online reviews that are spam or fake | -1.26 | -0.16 | Punishment | Yes |
| Education | College admissions decisions* | 0.40 | 1.13 | Reward | No |
| Education | Automated plagiarism detection | -0.16 | -1.47 | Punishment | No |
| Healthcare | Discharge from hospital* | 1.41 | 0.50 | Punishment | No |
| Business | Hiring based on resume screening | -0.18 | 1.04 | Reward | No |
| Business | Approving online reviews for publication | -1.15 | -0.06 | Reward | No |
| Banking | Detecting fraudulent credit card transactions | 0.44 | -1.19 | Reward | No |
| Banking | Approving loan applications | 0.49 | -0.01 | Reward | No |
| Government | Determining benefits eligibility* | 0.92 | -0.31 | Reward | No |
| Insurance | Adding an insurance surcharge if high-risk | 0.11 | -0.12 | Punishment | No |
| Insurance | Given an insurance discount if low-risk | -0.40 | -0.15 | Reward | No |
Appendix B Complete survey materials
B.1. Tasks
This section contains the full text used to describe each task at the beginning of Part 1.
Care plan
A hospital plans to use an AI system to identify patients who qualify for a preventative care plan, which includes access to educational materials and more frequent doctor check-ins. The AI will base its decisions on the patient’s health and past disease progression patterns. Doctors can override the AI’s decisions if needed. The hospital checks that the AI works well by seeing how often it agrees with doctors’ choices. The AI and the doctors agree most of the time.
Job probation
A company is planning to use an AI system to identify underperforming employees who need job probation. The AI will make its decisions based on performance metrics and work history. All flagged employees will be placed on probation, but they will not be demoted or fired unless a human review of their probationary period is negative. The company checks that the AI works well by seeing how often it agrees with human managers’ choices. The AI and the managers agree most of the time.
Tumor detection
A hospital plans to use AI system to look at CT scans to find potential cancerous tumors. If the AI detects signs of a tumor, the patient will receive additional tests. However, if no tumor is detected, additional tests will not be conducted. Doctors can override the AI’s decision if needed. The hospital checks that the AI works well by seeing how often it agrees with doctors’ choices. The AI and the doctors agree most of the time.
Tax audit
A government tax agency is planning to use an AI system to detect potential tax fraud. Everyone flagged by the AI will be audited, and audits will occur solely based on AI flags. However, taxpayers will only face penalties if a human investigation confirms the fraud. The tax agency checks that the AI works well by seeing how often it agrees with human experts’ choices of who to audit. The AI and the human experts agree most of the time.
Upgrade
An airline is planning to use an AI system to decide which passengers get upgrades to empty first-class seats. The AI will make its decisions based on factors like passenger loyalty status, ticket price, and check-in time. The AI’s decisions are final, but passengers can decline the upgrade and keep their original seat if they prefer. The airline checks that the AI works well by seeing how often it agrees with trained airline employees’ choices. The AI and the airline employees agree most of the time.
Rebook
An airline wants to use an AI system to manage overbooked flights. The AI will re-route passengers based on their destination, personal details (like age and group size), and booking details. While the AI’s decisions are final, passengers can opt for a refund or select an alternative re-routing plan if they are unhappy with the AI’s choice. The airline checks that the AI works well by seeing how often it agrees with trained airline employees’ choices. The AI and the employees agree most of the time.
App
A company is planning use an AI system to decide which customers get beta access to their new app, which offers exclusive deals and coupons. The AI will make its decisions based on expected financial gain for the company based on the customers’ past spending habits. The decisions about which customers get early access are final until the app is available to the public. The company checks that the AI works well by seeing how often it agrees with human marketing experts’ choices. The AI and the marketing experts agree most of the time.
Reviews
A website where people post reviews of businesses is planning to use an AI system to detect and remove fake reviews, e.g., posts by bots or by people who have not actually visited a business. Once a review is deleted, it cannot be recovered, but the person who wrote it can re-post a similar review if they want. The review website checks that the AI works well by seeing how often it agrees with trained human moderators’ choices. The AI and the moderators agree most of the time.
B.2. Multiplicity education
We provide the following text to participants at the beginning of the survey:
The image shows a table with 5 columns and two rows. The column headers are pictures, each of a cat or a dog. The two rows are labels AI#1 and AI#2. The table cells indicate what the AI models predict as the label for each image. Each cell includes the prediction (i.e., the word “cat” or the word “dog”) as well as a green checkmark, in the case of a correct prediction, or a red X in the case of an incorrect prediction. AI#1 is correct on 4/5 examples (it identifies one dog (column 3) as a cat). AI #2 is also correct on 4/5 examples, but identifies one cat (column 2) as a dog. Text to the right of each row points out that both AIs are correct on 4/5 examples.
When building Artificial Intelligence (AI), choosing the best model is crucial. AI developers consider a range of AI models and choose the one with the best performance. For example, if the goal is to distinguish between cats and dogs, developers will test different AIs using a set of images and choose the AI that correctly identifies the most pictures.
In the example below, AI #1 correctly identifies 4 out of 5 images. Suppose this is the highest-performing AI model available.
We display the top half of fig. 4 (i.e., images and AI #1 only).
Sometimes, however, multiple AI models perform equally well but differ in their specific errors. This is called model multiplicity. For example, AI #2 also correctly identifies 4 out of 5 images but makes a different mistake than AI #1.
We display fig. 4.
AI #1 and AI #2 are both well-suited for classifying cat and dog images. If no other AI performs better, developers must choose between AI #1 and AI #2.
Model multiplicity is common in AI: multiple models can perform equally well overall but make different mistakes. In some settings, like predicting cats and dogs, model multiplicity is inevitable but choosing between models is low-stakes. In other settings, choosing between models is more challenging, such as when the model predictions directly affect people.
Then, we ask the comprehension question “What does model multiplicity refer to?” with these choices: (a) It refers to a model making multiple correct predictions, (b) It means there are several models that perform equally well but makes different mistakes, (c) It describes an AI model that performs better than all other models, and (d) It refers to an AI model that can perform multiple task simultaneously.
B.3. Part 1: Fairness questions
We use two questions relating to distributive fairness, and two relating to procedural fairness. Each question is on a 7-point Likert scale (“Strongly disagree” to “Strongly agree”). The exact phrasing of each question is tailed to each scenario to be more descriptive (e.g., replacing the word “entity” with the context-appropriate word).
(Distributive) I think the AI will make good decisions.
(Distributive) I think the AI will make fair decisions.
(Procedural) I think it is appropriate for the entity to use AI in this context.
(Procedural) I would feel comfortable if the entity used AI to make a decision about me.
B.4. Part 1: Critical-thinking questions
We ask two critical-thinking questions midway through Part 1 of the survey to encourage participants to slow down and better absorb the material. The first question is multiple-choice, asking which of four statements is true: (a) The original AI model is the only good way to do [task], (b) The new AI and human experts always choose the same people, (c) Multiple models are equally good at [task], or (d) The original AI and new AI never agree on [task]. The correct answer is (c) and this is visible to participants as soon as they make any selection. The open-ended question asks “If the original AI and the new AI disagree about whether or not to [predict + or -], how should the [decision-making entity] decide what to do?” Respondents much enter at least 8 characters in their answer, but apart from that requirement there is no in-survey validation of the answer.
B.5. Part 2: Costs
Table 5 shows the (non-zero) cost that we use for each task.
| Task | Cost |
| Care plan | 15 minute longer wait time at appointments to see the doctor |
| Job probation | Employees will spend 15 extra minutes per week filling out paperwork about their work |
| Tumor detection | 15 minute longer wait time at appointments to see the doctor |
| Tax audit | There will be an additional $5 fee to file taxes |
| Upgrade | Airline tickets will cost $5 more |
| Rebook | Airline tickets will cost $5 more |
| App | The shipping time on all products will be 2 days longer |
| Reviews | The website will have to host more ads, including pop-up ads |
Appendix C Alternate statistical analyses
We describe two alternate statistical analyses for the Part 1 data (change in fairness perceptions after learning about multiplicity). As stated in section 4.4, we preregistered our study with a plan to use linear mixed models. Those results are in the main text. We include two alternate analyses here for two reasons: first, there is disagreement in the statistical community about what type of analysis is appropriate for Likert-scale data (Blanca Mena et al., 2017; Carifio and Perla, 2008; Jamieson, 2004; Kuzon et al., 1996; Norman, 2010). Including multiple options serves to “future-proof” our analysis in the case of future statistical consensus. Second, while preregistration is one tool to guard against “researcher degrees of freedom” (Simmons et al., 2011), another option is a multiverse analysis (Steegen et al., 2016). In a true multiverse analysis, researchers present the outcomes of all feasible statistical analyses on their data. Then, if the same conclusions hold for all or most of the analyses, they can be considered more robust. We present a partial multiverse analysis.
C.1. ART-ANOVA
Aligned-Rank-Transform (ART) ANOVA (Wobbrock et al., 2011) is a procedure that transforms ordinal (e.g., Likert-scale) data to fit the ANOVA assumptions. We fit our ART-ANOVA model and perform post-hoc contrast tests (Elkin et al., 2021) using the ARTool package in R (Kay et al., 2021). Our formula includes the factors time, task stakes (binary), task uncertainty (binary), and task framing and all possible interactions between factors. We also include the participant ID to account for within-subjects factor. We test H using a post-hoc contrast test with “time” as the factor of interest. We test H, H, and H with a post-hoc contrast test of the interaction term between “time” and the relevant task characteristic (with the “Interaction” flag set to true).
Results
Our ART-ANOVA model indicates that the following factors have significance: task stakes (F=35.2, Df=1, p), task framing (F=5.0, Df=1, p=0.026), and the interaction term between task stakes and task uncertainty (F=7.9, Df=1, p=0.005). As was the case when using LMMs in the analysis, none of the hypotheses H-H was significant.
C.2. Non-parametric tests
The Mann-Whitney U-test (Mann and Whitney, 1947) is another way we can evaluate H-H. This test is non-parametric, so does not require any assumptions about the data’s normality or distribution. We perform the tests using the wilcox.test function in R. For H, we test whether the average fairness perception is different as stratified by “time”. For H-H, we compute the change in score over time for each sample, and then conduct a Mann-Whitney U test on the data stratified by the relevant task characteristic.
Results
As is the case with using LMMs or ART-ANOVA, none of the hypotheses H-H is significant.
Appendix D Additional results (Fairness perceptions)
D.1. Task choice and baseline fairness perceptions
Table 6 shows the fairness perceptions for each task at the “pre” timestamp. These values should be interpreted as a baseline for how fair participants think it is to ML in each scenario. We observe that high-stakes tasks have lower fairness scores, as do punishment-framed tasks.
D.2. Demographic factors
We consider binary gender, integer age, binary experience with AI (defined as whether the participant currently uses AI in their work or school), and binary education (whether the participant has a bachelor’s degree or higher). There is no significant difference in initial fairness perceptions based on gender (t-test p=0.822), age (linear regression p=0.184), experience with AI (t-test p=0.088), or education (t-test p=0.328). There is no significant different in change in fairness scores after learning about multiplicity for gender (t-test p=0.921), age (linear regression p=0.944), AI experience (t-test p=0.460), or education (t-test p=0.413).
| Task | Score | Stakes | Unc. | Framing |
| Care plan | 0.26 (-0.75, 1.13) | High | High | Rew. |
| Job probation | -0.38 (-1.50, 0.88) | High | High | Pun. |
| Tumor detection | 0.12 (-0.75, 1.00) | High | Low | Rew. |
| Tax audit | 0.19 (-0.25, 1.00) | High | Low | Pun. |
| Upgrade | 1.01 ( 0.75, 1.75) | Low | High | Rew. |
| Rebook | 0.83 ( 0.50, 1.75) | Low | High | Pun. |
| App | 0.78 ( 0.00, 1.50) | Low | Low | Rew. |
| Reviews | 0.55 (-0.25, 1.50) | Low | Low | Pun. |
For high-stakes tasks, the mean fairness score is 0.05 on a 7-pt scale (IQR -0.81 to 1.0) and for low-stakes tasks it is 0.79 (IQR 0.25 to 1.5). This difference is statistically significant according to a Mann-Whitney U-test (W = 43476, p ). Punishment-framed tasks have a median fairness score of 0.30 (IQR -0.5 to 1.25), while for reward-framed tasks it is 0.55 (IQR 0.0 to 1.5). This difference is also statistically significant according to a Mann-Whitney U-test (W = 71016, p = 0.012). Task uncertainty does not have an impact on fairness score according to a Mann-Whitney U-test (W = 64358, p = 0.918).
Appendix E Additional results (Multiplicity resolution)
E.1. Details on testing H-H
Table 7 shows the effect sizes and significant for the post-hoc analyses conducted for H-H (section 5.2). We see that all effects except two are in the expected direction, and, using the Holm-Bonferroni correction, 10 of the 18 tests are significant.
E.2. Qualitative results
Table 8 shows the results of the open-ended question, where participants were asked how to resolve multiplicity prior to seeing the techniques that we proposed. The table contains 10 response categories and tallies the number of participants whose answer referenced that category for each of the 8 tasks. We did our best to interpret the responses, however, in some cases, we could not determine what the participant meant so we marked it as “unclear”. Most of these responses were on-topic and were a clear attempt to answer the question.
| Hypothesis | Factor | Diff | SE | t-stat | p-val | Sig? |
| H | human v. fancy | 0.188 | 0.025 | 7.563 | ¡0.0001 | Yes |
| H | human v. default-good | 0.214 | 0.026 | 8.382 | ¡0.0001 | Yes |
| H | human v. default-bad | 0.462 | 0.025 | 18.340 | ¡0.0001 | Yes |
| H | human v. ignore | 0.375 | 0.024 | 15.594 | ¡0.0001 | Yes |
| H | human v. random | 0.511 | 0.023 | 22.407 | ¡0.0001 | Yes |
| H | random v. fancy | -0.323 | 0.026 | -12.592 | ¡0.0001 | Yes |
| H | random v. default-good | -0.297 | 0.026 | -11.304 | ¡0.0001 | Yes |
| H | random v. default-bad | -0.049 | 0.026 | -1.887 | 0.0593 | No |
| H | random v. ignore | -0.136 | 0.025 | -5.477 | ¡0.0001 | Yes |
| H | human | 0.097 | 0.030 | 3.211 | 0.0013 | Yes |
| H | fancy | 0.043 | 0.038 | 1.128 | 0.2596 | No |
| H | ignore | -0.089 | 0.036 | -2.468 | 0.0137 | No |
| H | random | -0.079 | 0.033 | -2.411 | 0.0160 | No |
| H | human* | -0.004 | 0.031 | -0.133 | 0.8943 | No |
| H | default-good | 0.016 | 0.040 | 3.965 | ¡ 0.0001 | Yes |
| H | fancy | -0.059 | 0.039 | -1.510 | 0.1313 | No |
| H | random | 0.010 | 0.033 | 0.300 | 0.7641 | No |
| H | fancy* | 0.015 | 0.039 | 0.389 | 0.6970 | No |
| Care plan | Job | Tumor | Tax audit | Upgrade | Rebook | App | Reviews | Total | |
| probation | detection | ||||||||
| =43 | =43 | =45 | =45 | =46 | =45 | =46 | =45 | =358 | |
| human | 35 | 35 | 38 | 23 | 22 | 28 | 28 | 35 | 244 |
| default-good | 1 | 1 | 7 | 7 | 3 | 2 | 21 | ||
| ignore | 1 | 1 | 5 | 5 | 5 | 2 | 19 | ||
| fancy | 1 | 2 | 5 | 1 | 3 | 2 | 14 | ||
| random | 2 | 5 | 2 | 1 | 10 | ||||
| default-bad | 1 | 1 | 3 | 1 | 3 | 9 | |||
| Specific rule | 1 | 1 | 1 | 4 | 5 | 3 | 2 | 17 | |
| suggestion | |||||||||
| Ask the | 1 | 2 | 3 | 6 | |||||
| decision subject | |||||||||
| Unclear* | 2 | 3 | 2 | 7 | 4 | 6 | 2 | 5 | 31 |
| Not sure* | 1 | 2 | 1 | 4 |
E.3. Task-by-task preferences
Figure 5 shows the relative preferences for each multiplicity resolution technique for tasks stratified by stakes, uncertainty, and framing.
Marginal mean preferences over multiplicity resolution techniques stratified by task stakes, task framing, and task uncertainty. We report which results have non-overlapping confidence intervals for each setting. For task stakes, human (more preferred in high stakes), default bad (more preferred in low stakes), zero cost (more preferred in low stakes) and non-zero cost (more preferred in high stakes). For task framing, random (more preferred for reward tasks), default good (more preferred for reward), and default bad (more preferred for punishment). For task uncertainty, no factor has non-overlapping confidence intervals.
E.4. Costs
As noted in finding E4 in section 5.3, no extra cost is significantly preferred over extra cost in aggregate. However, for three high-stakes settings (tumor detection, care plan access, and job probation), this distinction is not significant. The point estimates for marginal mean preferences for extra cost are 0.436 (SE 0.032, p¡0.0001) for tumor detection, 0.474 (SE 0.038, p¡0.0001) for care plan access, and 0.445 (SE 0.035, p¡0.0001) for job probation. (A point estimate less than 0.50 indicates that no extra cost is more preferred, and a point estimate of 0.50 indicates that there is no preferences between the cost options.)
E.5. Demographic factors
There is no significant difference in preference for multiplicity resolution technique based on education level or AI experience. That is, a t-test comparing the results for each resolution technique has a p-value greater than 0.05. When we consider age as a binary factor (divided at 35, roughly the median age and a common cutoff for “young adult”), fancy is preferred by people younger than 35 (diff=0.081, SE=0.039, t= 2.054, p=0.040) and default_good is preferred by people older than 35 (diff=-0.101, SE=0.040, t=-2.515, p=0.012). Looking at gender, fancy is preferred by men (diff=0.0783, SE=0.039, t=2.009, p=0.045) and random is preferred by women (diff=-0.070, SE=0.035, t=-2.027, p=0.043). We encourage future work to explore whether these trends hold as the experiment setup is varied.